### v1

# Ground state solutions to the nonlinear Born-Infeld problem

Abstract.
In the paper we show the existence of ground state solutions to the nonlinear Born-Infeld problem
$\mathrm{div}\,\left(\frac{\nabla u}{\sqrt{1-|\nabla u|^{2}}}\right)+f(u)=0,\quad x\in\mathbb{R}^{N}$
in the zero and positive mass cases. Moreover, we find a new proof of the Sobolev-type inequality
$\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)\,dx\geq C_{N,p}\left(\int_{\mathbb{R}^{N}}|u|^{p}\,dx\right)^{\frac{N}{N+p}},$
for $p>2^{*}$ as well as the characterization of the optimal constant $C_{N,p}$ in terms of the ground state energy level. Previous approaches relied on approximation schemes and/or symmetry assumptions, which typically yield to compact embeddings and may lead to solutions that are not at the ground state energy level. In contrast, neither approximation arguments nor symmetry assumptions are employed in the paper to obtain a ground state solution. Instead, we develop a new direct variational approach based on minimization over a Pohožaev manifold combined with profile decomposition techniques. Finally, we show that nonradial solutions exist whenever $N\geq 4$; in particular, this settles a previously open problem in the case $N=5$.
Keywords: Born-Infeld equation, infinite energy problem, ground states, profile decompositions, nonradial solutions.
AMS Subject Classification: 35A15, 35J25, 35J93, 35Q75
1. Introduction
In this paper, we consider the nonlinear problem
(1.1) $\mathrm{div}\,\left(\frac{\nabla u}{\sqrt{1-|\nabla u|^{2}}}\right)+f(u)=0,\quad x\in\mathbb{R}^{N}.$
The problem (1.1) appears when one considers the electromagnetic theory introduced by Born and Infeld [13, 14, 15, 16] as a nonlinear alternative to the Maxwell theory. The new electromagnetic theory solves the well-known infinite energy problem. Namely, if we consider the electromagnetic field generated by a point charge in $\mathbb{R}^{3}$, and the corresponding scalar field equation in the Maxwell theory
$\Delta\phi+\delta_{0}=0,\quad x\in\mathbb{R}^{3},$
then the fundamental solution
$\phi(x)=\frac{C}{|x|}$
where $C>0$, has an infinite energy
$E_{M}(\phi)=\int_{\mathbb{R}^{3}}|\nabla\phi(x)|^{2}\,dx=+\infty.$
In the Born-Infeld theory the scalar field equation is governed by the operator $\mathrm{div}\,\left(\frac{\nabla\phi}{\sqrt{1-|\nabla\phi|^{2}}}\right)$. Then the fundamental solution $\phi(x)=\phi(|x|)$ to
$\mathrm{div}\,\left(\frac{\nabla\phi}{\sqrt{1-|\nabla\phi|^{2}}}\right)+\delta_{0}=0,\quad x\in\mathbb{R}^{3},$
is given by the derivative of its radial profile with some $b>0$
$\phi^{\prime}(r)=\frac{b^{2}}{\sqrt{r^{4}+b^{2}}}$
and its energy
$E_{BI}(\phi)=\int_{\mathbb{R}^{3}}\left(1-\sqrt{1-|\nabla\phi(x)|^{2}}\right)\,dx=4\pi\int_{0}^{\infty}\left(1-\frac{r^{2}}{\sqrt{r^{4}+b^{2}}}\right)r^{2}\,dr<\infty.$
We remark that (1.1) has variational structure. In fact, the energy functional associated to (1.1) is given by
(1.2) $I(u):=\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)\,dx-\int_{\mathbb{R}^{N}}F(u)\,dx=\Psi(u)-\Phi(u),$
where we set
$F(u):=\int_{0}^{u}f(s)\,ds,\quad\Psi(u):=\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)\,dx,\quad\Phi(u):=\int_{\mathbb{R}^{N}}F(u)\,dx.$
Formally, if $\|\nabla u\|_{\infty}=\|\nabla u\|_{L^{\infty}(\mathbb{R}^{N})}<1$, then
$\lim_{h\to 0}\frac{I(u+t\varphi)-I(u)}{h}=\int_{\mathbb{R}^{N}}\frac{\nabla u\cdot\nabla\varphi}{\sqrt{1-|\nabla u|^{2}}}-f(u)\varphi\,dx$
and $u$ is a weak solution of (1.1) provided $u$ is a critical point of $I$ with $\|\nabla u\|_{\infty}<1$. However, it should be pointed out that for the well-definedness of $\Psi$, the condition $|\nabla u|\leq 1$ for a.e. $x\in\mathbb{R}^{N}$ is required and it is not possible to find a function space on which $\Psi$ becomes of class ${\mathcal{C}}^{1}$. Therefore, we need to go beyond the smooth critical point theory to treat $I$ directly on some function space, and the nonsmoothness of $\Psi$ makes the problem complicated.
The problem
$\mathrm{div}\,\left(\frac{\nabla u}{\sqrt{1-|\nabla\phi|^{2}}}\right)+\rho=0$
with fixed $\rho$ has been studied by many authors; see, for example, [7, 8, 9, 10, 11, 17, 20, 25, 26]. On the other hand, the literature for (1.1) is fragmentary. In [12], the power nonlinearities $f(u)=|u|^{p-2}u$ with $p>2^{*}$ are studied and the existence of radial positive classical solution and infinitely many radial solutions is proved through approximations of the differential operator. Furthermore, the following Sobolev type inequality is established and the existence of optimizers is studied:
(1.3) $\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)dx\geq C_{N,p}\left(\int_{\mathbb{R}^{N}}|u|^{p}\,dx\right)^{\frac{N}{N+p}}.$
In [1], (1.1) is studied for $N\geq 2$ and f(u)=u+g(u)f(u)=-u+g(u) by the shooting method, and the existence of radial positive decreasing solution is found under some assumptions on $g$. On the other hand, [2] treats the case $N\geq 3$ and $f(u)=|u|^{p-2}u$ with $p>2$ and $p\neq 2^{*}$ where $2^{*}:=2N/(N-2)$. It is proved that for $p\in(2,2^{*})$, (1.1) has no radial solution decaying at infinity, and for $p>2^{*}$, (1.1) has infinitely many radial solutions with infinite energy, that is $\|\nabla u\|_{2}=\infty$. More general nonlinearities are considered in [30] for the zero and the positive mass case, and the existence of at least one radial solutions is established through the mountain pass theorem and a similar approximation in [12]. Finally, [18] develops the monotonicity trick due to [31, 24, 21] for nonsmooth functionals in the setting of [32] and the existence of infinitely many radial solutions of (1.1) is proved. In addition, the existence of nonradial solutions is also shown when $N=4$ or $N\geq 6$. We also refer to [8] for the regularity result of radial local minimizers and [3] for the $L^{2}$-normalized solutions. It is worth mentioning that in all these works, radial symmetry (or some symmetry of functions) plays a crucial role to overcome the lack of compactness. Due to the usage of the symmetry, the existence of nonradial solutions to (1.1) when $N=5$ is missing in [18].
In this paper, we aim to develop a concentration compactness argument or profile decompositions for the nonsmooth functional corresponding to (1.1) and this allows us to discuss the existence of a ground state solution to (1.1), i.e., a solution with the least possible energy, which seems to be the most relevant from the physical point of view. Moreover, the argument can be applied for the existence of nonradial solutions when $N=5$. For this purpose, we introduce a new direct variational approach based on minimization over a Pohožaev set, as we shall see later. Since we study (1.1) in the zero and the positive mass case as in [30, 18], two sets of assumptions are provided. After the existence of ground state solutions, nonradial solutions are studied. We start our considerations with the zero mass case in the spirit of Berestycki and Lions [5].
1.1. The zero mass case
In this subsection, the results for the zero mass case are stated. To explain our setting, we first introduce a function space we work on in this case. The following choice is taken from [17, 18]. See [8] for another choice of function space.
Let $N\geq 3$ and choose $j_{0}\in\mathbb{N}$ such that $\max\{2^{*},N\}<2j_{0}$. Then we introduce the function space
$X_{0}:=\left\{u\in L^{2^{*}}(\mathbb{R}^{N})\ :\ \|u\|_{X_{0}}:=\|\nabla u\|_{2}+\|\nabla u\|_{2j_{0}}<\infty\right\}.$
Thanks to $N<2j_{0}$, it is easily seen from Sobolev’s and the interpolation inequality that
(1.4) $X_{0}\subset W^{1,2^{*}}(\mathbb{R}^{N})\cap{\mathcal{C}}_{0}(\mathbb{R}^{N}),\quad{\mathcal{C}}_{0}(\mathbb{R}^{N}):=\left\{u\in{\mathcal{C}}(\mathbb{R}^{N})\ :\ \lim_{|x|\to\infty}u(x)=0\right\}.$
Note that if $u\in X_{0}$ and $\|\nabla u\|_{\infty}\leq 1$, then the inequality $1-\sqrt{1-t^{2}}\leq t^{2}$ in $[0,1]$ gives
$\Psi(u)=\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)\,dx\leq\int_{\mathbb{R}^{\mathbb{N}}}\left|\nabla u\right|^{2}dx<\infty.$
This observation allows us to define the domain of $\Psi$ as
$D(\Psi):=\left\{u\in X_{0}\ :\ \|\nabla u\|_{\infty}\leq 1\right\}.$
Then $\Psi:X_{0}\rightarrow\mathbb{R}\cup\{\infty\}$ is a convex functional with a convex domain $D(\Psi)$.
Next, to treat $\Phi$ in (1.2) on $X_{0}$, we introduce the following assumptions on $f$:
  1. (F1)
    $f\in{\mathcal{C}}^{1}(\mathbb{R})$ and $f(0)=0$;
  2. (F2)
    $\limsup_{s\to 0}\frac{f^{\prime}(s)}{|s|^{2^{*}-2}}=0$;
  3. (F3)
    $sH^{\prime}(s)>0$ for all $s\neq 0$ where $H(s):=F(s)+\frac{1}{N}f(s)s$;
  4. (F4)
    $\lim_{|s|\to\infty}H(s)=\infty$;
  5. (F5)
    $s\mapsto f(s)s$ is nonincreasing in $(-\infty,0)$ and nondecreasing in $(0,\infty)$;
  6. (F6)
    $|F(s)|\lesssim|f(s)s|$ for $s\in\mathbb{R}$, that is, there exists a $C\geq 0$ such that $|F(s)|\leq C|f(s)s|$ for every $s\in\mathbb{R}$.
Remark 1.1.
  1. (i)
    Under (F1), condition (F5) is equivalent to $0\leq f^{\prime}(s)s^{2}+f(s)s$ for every $s\in\mathbb{R}$. Thus, a direct computation shows that (F1) and (F5) with an additional condition $sf(s)>0$ for each $s\neq 0$ imply (F3).
  2. (ii)
    Condition (F6) cannot be derived from (F1)–(F5) and such an example of $f$ is given as follows. Let $k_{0}\in\mathbb{N}$ satisfy $2^{*}<2k_{0}$ and consider $F(s):=\log(1+s^{2k_{0}})\in{\mathcal{C}}^{\infty}(\mathbb{R})$. Since $f(s)=2k_{0}s^{2k_{0}-1}/(1+s^{2k_{0}})$, it is immediate to verify (F1)–(F5), however, (F6) is not satisfied.
By (1.4), (F1) and (F2), we may verify $\Phi\in{\mathcal{C}}^{1}(X_{0};\mathbb{R})$ and it allows us to define the notion of a critical point of $I$ as in [32]. Namely, $u\in X_{0}$ is said to be a critical point of $I$ if
(1.5) $\Psi(v)-\Psi(u)-\Phi^{\prime}(u)(v-u)\geq 0\quad\text{for all $v\in X_{0}$},$
that is, $\Phi^{\prime}(u)\in\partial\Psi(u)$, where $\partial\Psi(u)$ denotes the subdifferential of a convex functional $\Psi$.
To discuss the existence of ground state solutions of (1.1), we will introduce the Pohožaev set ${\mathcal{M}}_{0}$ for (1.1). Let
(1.6) $M_{0}(u):=\Psi(u)-\int_{\mathbb{R}^{N}}H(u)\,dx=\Psi(u)-\Phi(u)-\frac{1}{N}\Phi^{\prime}(u)(u)$
and define ${\mathcal{M}}_{0}$ by
(1.7) ${\mathcal{M}}_{0}:=\left\{u\in X_{0}\setminus\{0\}\ :\ M_{0}(u)=0\right\}.$
Since for $u\in X_{0}\setminus D(\Psi)$ $\Psi(u)=\infty$ and $M_{0}(u)=\infty$ hold, we easily get ${\mathcal{M}}_{0}\subset D(\Psi)\setminus\{0\}$. For the semilinear scalar field problems, it is known that each reasonable solution satisfies the Pohožaev identity (for instance, see [5, Proposition 1]), and hence any reasonable solution belongs to the Pohožaev set. A counterpart for (1.1) of this fact is obtained in [18, Proposition 2.6] (see also Section˜2). Therefore, we are led to the following minimizing problem:
$c_{0}:=\inf_{{\mathcal{M}}_{0}}I.$
Remark that if there exists a minimizer $u$ corresponding to $c_{0}$ and if $u$ is a nontrivial solution of (1.1), then $u$ has the least energy among all nontrivial solutions of (1.1) in $X_{0}$, and hence $u$ is a ground state solution of (1.1).
Theorem 1.2.
Let $N\geq 3$ and assume (F1)–(F6). Then the following assertions hold:
  1. (i)
    $0<c_{0}=c_{0,MP}$ holds where $c_{0,MP}$ is a mountain pass level given by (4.3);
  2. (ii)
    every minimizing sequence $(u_{n})_{n\in\mathbb{N}}$ for $c_{0}$ is relatively compact in $X_{0}$ up to translations, and if $\|u_{n}-u_{0}\|_{X_{0}}\to 0$, then $u_{0}\in{\mathcal{M}}_{0}$ attains $c_{0}$ and is a ground state solution of (1.1), namely $u_{0}$ is a classical solution of (1.1) with $I(u_{0})=c_{0}$;
  3. (iii)
    for each ground state solution $u$ of (1.1), either $u>0$ in $\mathbb{R}^{N}$ or $u<0$ in $\mathbb{R}^{N}$ holds, and there exists a ground state solution of (1.1) which is radially symmetric.
Remark 1.3.
By Theorem˜1.2 (iii), the ground state energy of (1.1) is equal to the radial ground state energy, that is, the least energy among all radial nontrivial solutions of (1.1) in $X_{0}$. However, a priori, it is not clear whether the ground state energy coincides with the radial one since we do not impose on $f$ the oddness.
When $f(u)=|u|^{p-2}u$ with $p>2^{*}$, as in the seminlinear case (see [29]), the existence of minimizers corresponding to $c_{0}$ implies the Sobolev type inequality (1.3) obtained in [12], the expression of the best constant with $c_{0}$ and the characterization of every optimizer with the ground state solutions of (1.1):
Theorem 1.4.
Let $N\geq 3$ and $p>2^{*}$. Then for any $u\in X_{0}$ there holds
(1.8) $\int_{\mathbb{R}^{N}}\left(1-\sqrt{1-|\nabla u|^{2}}\right)\,dx\geq\left(\frac{N+p}{Np}\right)N^{\frac{p}{N+p}}\left(\inf_{{\mathcal{M}}_{0}}I\right)^{\frac{p}{N+p}}\left(\int_{\mathbb{R}^{N}}|u|^{p}\,dx\right)^{\frac{N}{N+p}}.$
Moreover, every optimizer $v$ has a form of $v(x)=\theta u(x/\theta)$, where $\theta>0$ and $u$ is a ground state solution of (1.1) with $f(u)=|u|^{p-2}u$.
Remark 1.5.
  1. (1)
    A difference between [12, Theorem 1.2] and Theorem˜1.4 is in the characterization of optimizer through ground state solutions of (1.1) with $f(u)=|u|^{p-2}u$. In fact, in [12, Theorem 1.2] it is shown that the constant is achieved by a radial solution of (1.1).
  2. (2)
    When $p=2^{*}$, we may also obtain a similar inequality to (1.8) with the best constant involving $\inf_{{\mathcal{M}}_{0}}I$. However, in this case, the constant is never achieved as shown in [12].
1.2. Positive mass case
In this case, let $N\geq 1$ and suppose that the nonlinearity $f$ is of the form f(s)=s+g(s)f(s)=-s+g(s), where $g\in{\mathcal{C}}^{1}(\mathbb{R})$ satisfies the following conditions:
  1. (G1)
    $g^{\prime}(0)=0=g(0)$;
  2. (G2)
    the map $s\mapsto\frac{g(s)}{|s|}$ is strictly increasing on $\mathbb{R}$;
  3. (G3)
    $\lim_{|s|\to\infty}\frac{g(s)}{s}=\infty$.
In this case, choose $j_{0}$ such that $2j_{0}>N$ and introduce the space
$X_{1}:=\{u\in H^{1}(\mathbb{R}^{N})\ :\ \nabla u\in L^{2j_{0}}(\mathbb{R}^{N})\}$
with the norm
$\|u\|_{X_{1}}:=\|u\|_{H^{1}}+\|\nabla u\|_{2j_{0}}.$
Then, the variational functional $I$ is given by (1.2), $\Psi$ is defined as before with the natural domain
$D(\Psi):=\left\{u\in X_{1}\ :\ \|\nabla u\|_{\infty}\leq 1\right\}$
and
$\Phi(u):=\int_{\mathbb{R}^{N}}-\frac{1}{2}u^{2}+G(u)\,dx,\quad G(s):=\int_{0}^{s}g(t)\,dt.$
Remark that $\Psi$ is a convex functional on a convex domain $D(\Psi)$ and $\Phi\in{\mathcal{C}}^{1}(X_{1};\mathbb{R})$. Therefore, the notation of critical points of $I$ on $X_{1}$ is defined as in (1.5) by changing $X_{0}$ to $X_{1}$. We also introduce the Pohožaev set as
${\mathcal{M}}_{1}:=\left\{u\in X_{1}\setminus\{0\}\ :\ M_{1}(u)=0\right\}\subset D(\Psi)\setminus\{0\},\quad M_{1}(u):=\Psi(u)-\Phi(u)-\frac{1}{N}\Phi^{\prime}(u)(u)$
and consider the following minimizing problem:
$c_{1}:=\inf_{{\mathcal{M}}_{1}}I.$
Then $c_{1}$ gives the ground state energy as follows:
Theorem 1.6.
Suppose $N\geq 1$, f(s)=s+g(s)f(s)=-s+g(s) and (G1)–(G3). Then
  1. (i)
    $0<c_{1}=c_{1,MP}$ holds where $c_{1,MP}$ is a mountain pass level;
  2. (ii)
    every minimizing sequence $(u_{n})_{n\in\mathbb{N}}$ for $c_{0}$ is relatively compact in $X_{1}$ up to translations, and if $\|u_{n}-u_{0}\|_{X_{1}}\to 0$, then $u_{1}\in{\mathcal{M}}_{1}$ attains $c_{1}$ and is a ground state solution of (1.1), namely $u_{0}$ is a classical solution of (1.1) with $I(u_{0})=c_{1}$;
  3. (iii)
    for each ground state solution $u$ of (1.1), either $u>0$ in $\mathbb{R}^{N}$ or $u<0$ in $\mathbb{R}^{N}$ holds, and there exists a ground state solution of (1.1) which is radially symmetric.
1.3. Nonradial solutions
Now we show that (1.1) also admits nonradial solutions provided $N\geq 4$. Let $X$ denote the space $X_{0}$ in the zero-mass case, or $X_{1}$ in the positive-mass case. Similarly, let $\mathcal{M}$} denote the corresponding Pohožaev set, ${\mathcal{M}}_{0}$ or ${\mathcal{M}}_{1}$, respectively. We treat both cases simultaneously and assume either (F1)–(F6) or (G1)–(G3), depending on which case is considered.
The approach here is taken from [4, 6, 22, 28, 29, 34]. Let $k_{1},k_{2}\in\mathbb{N}$ satisfy $k_{1},k_{2}\geq 2$ and $N-k_{1}-k_{2}\geq 0$ and write $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\mathbb{R}^{N-k_{1}-k_{2}}$ (if $N=k_{1}+k_{2}$, then we write $x=(x_{1},x_{2})\in\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}$). Define
${\mathcal{O}}$ $={\mathcal{O}}(k_{1})\times{\mathcal{O}}(k_{2})\times\{\mathrm{id}\}\subset{\mathcal{O}}(N),$
$X_{{\mathcal{O}}}$ $=\left\{u\in X\ :\ u(\rho x)=u(x)\quad\text{for any $x\in\mathbb{R}^{N}$ and $\rho\in{\mathcal{O}}$}\right\}.$
For each $u\in X_{\mathcal{O}}$, we identity $u(x_{1},x_{2},x_{3})$ with $u(r_{1},r_{2},x_{3})$ where $r_{1}=|x_{1}|$ and $r_{2}=|x_{2}|$. Then we define $\tau$ by
$\tau(r_{1},r_{2},x_{3}):=(r_{2},r_{1},x_{3}),\quad(\tau u)(x_{1},x_{2},x_{3}):=-u(\tau(|x_{1}|,|x_{2}|,x_{3}))=-u(|x_{2}|,|x_{1}|,x_{3}).$
Remark that $\tau:X_{\mathcal{O}}\to X_{\mathcal{O}}$ is an isometry. We also set
$X_{\tau}:=\left\{u\in X_{\mathcal{O}}\ :\ \tau u=u\right\}.$
Clearly, if $u\in X_{\tau}$ is radial, i.e. $u(x)=u(\rho x)$ for any $\rho\in{\mathcal{O}}(N)$, then $u\equiv 0$. Hence $X_{\tau}$ does not contain any nontrivial radial functions.
Theorem 1.7.
Let $N\geq 4$ and fix $k_{1},k_{2}\in\mathbb{N}$ so that $k_{1},k_{2}\geq 2$ and $N-k_{1}-k_{2}\geq 0$. Suppose (F1)–(F6) (resp. (G1)–(G3)) and that $f$ (reps. $g$) is odd. Then there is a nontrivial critical point $u_{0}\in{\mathcal{M}}\cap X_{\tau}=:{\mathcal{M}}_{\tau}$ such that $I(u_{0})=\inf_{{\mathcal{M}}_{\tau}}I\geq\inf_{{\mathcal{M}}}I>0$. Moreover, when $k_{1}=k_{2}$, then
$I(u_{0})=\inf_{{\mathcal{M}}_{\tau}}I>2\inf_{{\mathcal{M}}}I>0.$
In particular $u_{0}$ is a nonradial classical solution of (1.1).
1.4. Comments
We describe arguments to prove Theorems˜1.2 and 1.6 and their difficulties. As explained above, in the literature, the existence of ground state solutions of (1.1) with a general nonlinearity $f$ is not studied. As pointed as in Subsection˜1.1, it is not clear that radial functions provide the ground state energy of (1.1) a priori, and hence the radial symmetry may not be exploited to prove the existence of ground state solutions for a general nonlinearity $f$. Thus, we face the lack of compact embedding of $X_{0}$ or $X_{1}$, and to the best of the authors’ knowledge, this paper is the first to deal with the lack of compactness for (1.1). In addition to the lack of compact embedding, the nonsmoothness of functionals $I,M_{0},M_{1}$ also makes the analysis complicated and different from the semilinear case (for instance, [5, 28, 29]). Thus, to solve difficulties and to prove Theorems˜1.2 and 1.6, new ideas or arguments are necessary.
To show the existence of minimizers corresponding to $c_{0}$ and $c_{1}$, consider any minimizing sequence $(u_{n})_{n\in\mathbb{N}}\subset X_{0}$ for $c_{0}$ (resp. $(u_{n})_{n\in\mathbb{N}}\subset X_{1}$ for $c_{1}$). Since the Pohožaev functionals $M_{0}$ and $M_{1}$ are different from the semilinear case, the boundedness of $(u_{n})_{n\in\mathbb{N}}$ in $X_{0}$ (resp. $X_{1}$) is an issue. To overcome this point, in the zero mass case, (F6) is used. On the other hand, in the positive mass case, the case $N\geq 3$ is simple and similar to the semilinear case, however, the cases $N=1,2$ are different. See the beginning of the proof of Theorem˜1.6.
After establishing the boundedness of $(u_{n})_{n\in\mathbb{N}}$, the next issue is to prove the relative compactness of $(u_{n})_{n\in\mathbb{N}}$ up to translations. Here we employ and develop profile decompositions suited for the nonsmooth functionals $M_{0},M_{1}$ and smooth functional $\Phi$, and our argument is inspired by [29, 19]. For the existence of minimizers, the projection to the Pohožaev sets ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$ are constructed and we compute and compare energies of projected profiles of $(u_{n})_{n\in\mathbb{N}}$ with $c_{0}$ and $c_{1}$. This detailed analysis also allows us to prove that every minimizer has a constant sign, and this fact leads to the existence of radial minimizers.
Even though we know the existence of minimizers, it is worth noting that whether these minimizers become critical points of $I$ is not trivial or immediate, and this point is more delicate than the semilinear case. For the semilinear case, a similar assertion is proved in [29, Lemma 3.3] where the differentiability of functional originated from the differential operator is utilized. In our case, the corresponding functional is $\Psi$, which is not smooth on $X_{0}$ or $X_{1}$. In addition, the nonsmoothness of $M_{0}$ and $M_{1}$ affect the properties of the projections to ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$, and the analysis of the projections onto ${\mathcal{M}}_{0}$, ${\mathcal{M}}_{1}$ and functional $I$ is more delicate and complicated. These issues will be solved in Subsections˜4.3 and 5.2 and it is one of novelties in this paper.
We emphasize that since the method in this paper does not rely on the compactness of the embedding of function spaces, our argument can be applied to obtain the existence of nonradial solutions to (1.1). In particular, Theorem˜1.7 handles the case $N=5$ and extends the results in [18]. Our setting is slightly different from [4, 6, 22, 28, 29, 34], and any combination of $k_{1},k_{2}$ with $k_{1},k_{2}\geq 2$ and $N-k_{1}-k_{2}\geq 0$ can be treated and we obtain other type of nonradial solutions which are not treated in [18].
It is also worth mentioning that our argument may be extended to a more general problem, for instance $f(x,u)$ in place of $f(u)$ in (1.1). In fact, for the semilinear case, the characterization of the ground state energy by the mountain pass level is obtained in [23], and it is useful to handle the nonlinearity $f(x,u)$ or a potential term $V(x)u$. The same characterization for (1.1) is obtained in Theorems˜1.2 and 1.6, and we expect that this might be helpful to study (1.1) with $f(x,u)$ or $V(x)u$. Furthermore, though it is only for minimizers, we succeed to prove that the constraint $M_{0}(u)=0$ (resp. $M_{1}(u)=0$) is natural in the sense that a critical point of $I|_{{\mathcal{M}}_{0}}$ (resp. $I|_{{\mathcal{M}}_{1}}$) provides a critical point of $I$. Therefore, it is interesting to see whether the same is true for higher energies or other nonsmooth constraints.
The structure of the paper is as follows. In Section˜2, the functional setting is recalled and some preliminary properties related to the functionals $I$ and $\Psi$ are proved. In Section˜3 we develop profile decompositions for bounded sequences in $X_{0}$ and $X_{1}$. Section˜4 is then devoted to the zero mass case and contains the proofs of Theorems˜1.2 and 1.4. In Section˜5 we turn to the positive mass case and prove Theorem˜1.6. Finally, in Section˜6 we establish the existence of nonradial solutions.
2. Concentration-compactness and a profile decomposition
In this section, preliminary results are stated and proved in the zero mass and positive mass cases. Throughout this section and Section˜3, we assume the following:
[…]
and set
$\|u\|_{X}:=\begin{dcases}\|u\|_{X_{0}}&\text{(the zero mass case)},\\ \|u\|_{X_{1}}&\text{(the positive mass case)}.\end{dcases}$
Then $(X,\|\cdot\|_{X})$ is a reflexive Banach space (cf. [17, Proposition 3.3]) and the following embedding holds (see (1.4) for the definition of ${\mathcal{C}}_{0}(\mathbb{R}^{N})$):
(2.2) $X\subset W^{1,p_{0}}(\mathbb{R}^{N})\cap{\mathcal{C}}_{0}(\mathbb{R}^{N}).$
Remark also that $D(\Psi)=\{u\in X\ :\ \|\nabla u\|_{\infty}\leq 1\}$ is convex in $X$. Similarly, put
$M(u):=\begin{dcases}M_{0}(u)&\text{(the zero mass case)},\\ M_{1}(u)&\text{(the positive mass case)}.\end{dcases}$
Finally, the space dimension $N$ is assumed to be $N\geq 3$ in the zero mass case and $N\geq 1$ in the positive mass case.
We first prove that the functional $\Psi$ restricted on $D(\Psi)$ is continuous:
Lemma 2.1.
Assume $(u_{n})_{n\in\mathbb{N}}\subset D(\Psi)$, $u\in D(\Psi)$ and $\|u_{n}-u\|_{X}\to 0$ as $n\to\infty$. Then $\Psi(u_{n})\to\Psi(u)$ as $n\to\infty$.
Proof.
Notice that
(2.3) $\frac{1}{2}t^{2}\leq\sqrt{1-t^{2}}-1\leq t^{2}\quad\text{for any $t\in[-1,1]$}.$
Remark also that $\|u_{n}-u\|_{X}\to 0$ yields $\|\nabla u_{n}-\nabla u\|_{2}\to 0$. Then we may find a subsequence $(u_{n_{k}})_{k\in\mathbb{N}}$ and $w\in L^{2}(\mathbb{R}^{N})$ such that $\nabla u_{n_{k}}(x)\to\nabla u(x)$ a.e. $\mathbb{R}^{N}$ and $|\nabla u_{n_{k}}(x)|+|\nabla u(x)|\leq w(x)$ a.e. $\mathbb{R}^{N}$. By (2.3) with $|\nabla u(x)|\leq 1$ and $|\nabla u_{n_{k}}(x)|\leq 1$ due to $u_{n_{k}},u\in D(\Psi)$, the dominated convergence theorem gives $\Psi(u_{n_{k}})\to\Psi(u)$ as $k\to\infty$. Since the limit is independent of subsequences, we have $\Psi(u_{n})\to\Psi(u)$ as $n\to\infty$ and this completes the proof. ∎
Next we introduce
$\Psi_{0}(u):=\begin{dcases}\Psi(u)&\text{(the zero mass case)},\\ \Psi(u)+a\|u\|^{2}_{2}&\text{(the positive mass case)},\end{dcases}$
where $a\in(0,\infty)$ is any fixed number.
Lemma 2.2.
Suppose that $(u_{n})_{n\in\mathbb{N}}\subset D(\Psi)$ satisfies $u_{n}\rightharpoonup u$ weakly in $X$ and $\Psi_{0}(u_{n})\to\Psi_{0}(u)$. Then $\|u_{n}-u\|_{X}\to 0$ holds.
Proof.
[…]
$\Psi_{0}(u)=\Psi(u)+a\|u\|_{2}^{2}<\lim_{k\to\infty}\Psi(u_{n_{k}})+a\liminf_{k\to\infty}\|u_{n_{k}}\|_{2}^{2}=\liminf_{k\to\infty}\Psi_{0}(u_{n_{k}})=\Psi_{0}(u),$
which is a contradiction. Hence $\limsup_{n\to\infty}\Psi(u_{n})\leq\Psi(u)$ and $\Psi(u_{n})\to\Psi(u)$ holds. In a similarly way, $\|u_{n}\|_{2}\to\|u\|_{2}$ can be proved and $\|u_{n}-u\|_{2}\to 0$ holds in the positive mass case.
To prove $\|\nabla u_{n}\|_{2j}\to\|\nabla u\|_{2j}$ from $\Psi(u_{n})\to\Psi(u)$, let us recall the following expansion ([25])
(2.4) $1-\sqrt{1-t^{2}}=\sum_{j\in\mathbb{N}}b_{j}t^{2j},\quad b_{j}>0.$
In particular,
$\Psi(u)=\sum_{j\in\mathbb{N}}b_{j}\|\nabla u\|_{2j}^{2j}\quad\text{for any $u\in D(\Psi)$}.$
For each $k\in\mathbb{N}$, set
$\Psi_{k,1}(u):=\sum_{j=1}^{k}b_{j}\|\nabla u\|_{2j}^{2j},\quad\Psi_{k,2}(u):=\sum_{j=k+1}^{\infty}b_{j}\|\nabla u\|_{2j}^{2j}.$
It is not hard to see that $\Psi_{k,1}$ and $\Psi_{k,2}$ are convex on $D(\Psi)$, and weakly lower semicontinuous. By utilizing the above argument and $\Psi(u_{n})\to\Psi(u)$, we obtain $\Psi_{k,1}(u_{n})\to\Psi(u)$ and $\Psi_{k,2}(u_{n})\to\Psi(u)$ for each $k\in\mathbb{N}$. By starting from the case $k=1$, $\|\nabla u_{n}\|_{2j}^{2j}\to\|\nabla u\|_{2j}^{2j}$ is obtained for every $j\geq 1$. This completes the proof. ∎
Finally, we prove that every critical point of $I$ in $X$ becomes a classical solution of (1.1) and satisfies the Pohožaev identity. A similar assertion is obtained in [18, Propositions 2.6 and 2.10] for $N\geq 3$. Here our main focus is in the case $N=1$. Notice that under (F1)–(F2) (resp. (G1)), $\Phi\in{\mathcal{C}}^{1}(X;\mathbb{R})$ due to (2.2).
Lemma 2.3.
Suppose either $N\geq 3$ and (F1)–(F2) or else $N\geq 1$ and (G1). Then each critical point $u$ of $I$ in $X$ is a classical solution of (1.1) and the Pohožaev identity $M(u)=0$ holds.
Proof.
Let $u\in X$ be a critical point of $I$ and define
$\rho(x):=\begin{dcases}f(u(x))&\text{(the zero mass case)},\\ -u(x)+g(u(x))&\text{(the positive mass case)}.\end{dcases}$
By (2.2), $\rho\in L^{\infty}(\mathbb{R}^{N})\cap{\mathcal{C}}(\mathbb{R}^{N})$ holds and by rewriting (1.5), $u$ is a (unique) minimizer of
$J(v):=\Psi(v)-\int_{\mathbb{R}}\rho v\,dx:X\to\mathbb{R}\cup\{\infty\}.$
When $N\geq 3$, from [18, Propositions 2.13] (cf. [18, Proposition 2.10]), we see that $u$ is a weak solution of (1.1) and $\|\nabla u\|_{\infty}<1$ holds. In particular, by [11] (cf. [17, Proof of Theorem 1.11 Step 5]), $u\in{\mathcal{C}}^{1,\alpha}(\mathbb{R}^{N})$ for some $\alpha\in(0,1)$ and $u$ is a strong solution of
(2.5) $\sum_{i,j=1}^{N}a_{ij}(x)\partial_{i}\partial_{j}u=\rho\quad\text{in}\ \mathbb{R}^{N},\quad a_{ij}(x):=\frac{\delta_{ij}}{\sqrt{1-|\nabla u(x)|^{2}}}+\frac{\partial_{i}u(x)\partial_{j}u(x)}{\left(1-|\nabla u(x)|^{2}\right)^{3/2}}$
Since $a_{ij}\in{\mathcal{C}}^{0,\alpha}_{\rm loc}(\mathbb{R}^{N})$ and $\rho\in{\mathcal{C}}^{1}(\mathbb{R}^{N})$, the Schauder theory gives $u\in{\mathcal{C}}^{2}(\mathbb{R}^{N})$.
We remark that the argument in [18] is also valid for $N=2$ in the positive mass case. Hence, in this case, $u\in{\mathcal{C}}^{1,\alpha}(\mathbb{R}^{2})$ is a strong solution of (2.5) and the argument above leads to $u\in{\mathcal{C}}^{2}(\mathbb{R}^{2})$.
Let $N=1$ and $u\in X$ be a critical point of $I$, and notice that $\rho\in L^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ in view of (G1) and (2.2). We follow the argument in [8, Theorem 1.4, Proposition 2.7, Remark 2.9]. Without any changes, the argument in [8, Proposition 2.7] shows that
(2.6) $\frac{(u^{\prime})^{2}}{\sqrt{1-(u^{\prime})^{2}}}\in L^{1}(\mathbb{R}),\quad\frac{u^{\prime}\psi^{\prime}}{\sqrt{1-(u^{\prime})^{2}}}\in L^{1}(\mathbb{R})\quad\text{for each $\psi\in X_{1}$ with $\psi^{\prime}\in L^{\infty}(\mathbb{R})$}.$
We next prove that $u$ is a weak solution of (1.1) by a slightly modified argument in [8, Proof of Theorem 1.4]. Let $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$ and $E_{k}:=\{x\in\mathbb{R}\ |\ |u^{\prime}(x)|\geq 1-1/k\}$ for $k\in\mathbb{N}$.
If $|E_{k}|=0$ holds for some $k\in\mathbb{N}$, then $\|u^{\prime}\|_{\infty}<1$ and $u+t\psi\in D(\Psi)$ provided $|t|\ll 1$. Since $u$ is a minimizer of $J$, $J(u)\leq J(u+t\psi)$ holds for any $t\in\mathbb{R}$ with $|t|\ll 1$. This together with $\|u^{\prime}\|_{\infty}<1$ and the dominated convergence theorem implies
$\int_{\mathbb{R}}\frac{u^{\prime}\psi^{\prime}}{\sqrt{1-(u^{\prime})^{2}}}\,dx=\int_{\mathbb{R}}\rho\psi\,dx=\int_{\mathbb{R}}f(u)\psi\,dx\quad\text{for all $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$}.$
In particular, $u$ is a weak solution of (1.1). Furthermore, by $\|u^{\prime}\|_{\infty}<1$, the above equality asserts that $u^{\prime}/\sqrt{1-(u^{\prime})^{2}}\in W^{1,\infty}(\mathbb{R})$ with the derivative $\rho\in{\mathcal{C}}^{1}(\mathbb{R})$. Thus, $u\in{\mathcal{C}}^{2}(\mathbb{R})$.
In what follows, we may suppose $|E_{k}|>0$ for each $k\in\mathbb{N}$. Notice that $E_{k+1}\subset E_{k}$, $|E_{2}|<\infty$ and $|E_{k}|\to 0$ as $k\to\infty$ by (2.6). Let $n_{0}\in\mathbb{N}$ be large enough so that $0<|[-n_{0},n_{0}]\cap E_{2}^{c}|$ and set
$\eta_{0}(x):=\frac{1}{|[-n_{0},n_{0}]\cap E_{2}^{c}|}\int_{-\infty}^{x}\bm{1}_{[-n_{0},n_{0}]\cap E_{2}^{c}}(s)\,ds\in W^{1,\infty}(\mathbb{R}).$
For $k\geq 2$, introduce
$a_{k}:=\int_{\mathbb{R}}\psi^{\prime}(s)\bm{1}_{E_{k}^{c}}(s)\,ds=\int_{E_{k}^{c}}\psi^{\prime}(s)\,ds\in\mathbb{R}.$
Since $|E_{k}|\to 0$ as $k\to\infty$, the fact $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$ yields
$\lim_{k\to\infty}a_{k}=\int_{\mathbb{R}}\psi^{\prime}\,ds=0.$
Define $\psi_{k}$ by
$\psi_{k}(x):=\int_{-\infty}^{x}\psi^{\prime}(s)\bm{1}_{E_{k}^{c}}(s)\,ds-a_{k}\eta_{0}(x).$
Let $R>n_{0}$ satisfy $\mathrm{supp}\,\psi\subset[-R,R]$. Since $\eta_{0}$ and $\psi_{k}$ are constant on $\mathbb{R}\setminus[-R,R]$, the choice of $a_{k}$ yields $\mathrm{supp}\,\psi_{k}\subset[-R,R]$ for every $k\geq 2$ and $|\psi_{k}^{\prime}(x)|>0$ implies $x\in E_{k}^{c}$. Furthermore, from $a_{k}\to 0$ and $|E_{k}|\to 0$ with $\mathrm{supp}\,\psi_{k},\mathrm{supp}\,\psi\subset[-R,R]$, it follows that $\|\psi_{k}-\psi\|_{1}\to 0$. By $\|u^{\prime}+t\psi_{k}^{\prime}\|_{\infty}\leq 1-1/(2k)$ for $|t|\ll 1$, the dominated convergence theorem yields
$\lim_{t\to 0}\frac{J(u+t\psi_{k})-J(u)}{t}$ $=\int_{\mathbb{R}}\frac{u^{\prime}\psi_{k}^{\prime}}{\sqrt{1-(u^{\prime})^{2}}}\,dx-\int_{\mathbb{R}}\rho\psi_{k}\,dx$
$=\int_{\mathbb{R}}\frac{u^{\prime}\left(\psi^{\prime}\bm{1}_{E_{k}^{c}}-a_{k}\eta_{0}^{\prime}\right)}{\sqrt{1-(u^{\prime})^{2}}}\,dx-\int_{\mathbb{R}}\rho\psi_{k}\,dx.$
Recalling that $u$ is a minimizer of $J$, we observe that for each $k\geq 2$ and $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$,
$0\leq\lim_{t\to 0}\frac{J(u+t\psi_{k})-J(u)}{t}=\int_{\mathbb{R}}\frac{u^{\prime}\left(\psi^{\prime}\bm{1}_{E_{k}^{c}}-a_{k}\eta_{0}^{\prime}\right)}{\sqrt{1-(u^{\prime})^{2}}}\,dx-\int_{\mathbb{R}}\rho\psi_{k}\,dx.$
Letting $k\to\infty$ together with (2.6) and $\|\psi_{k}-\psi\|_{1}\to 0$ leads to
$0\leq\int_{\mathbb{R}}\frac{u^{\prime}\psi^{\prime}}{\sqrt{1-(u^{\prime})^{2}}}\,dx-\int_{\mathbb{R}}\rho\psi\,dx\quad\text{for every $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$}.$
Since $\psi\in{\mathcal{C}}^{\infty}_{c}(\mathbb{R})$ is arbitrary, the above inequality becomes the equality and $u$ is a weak solution of (1.1), $u^{\prime}/\sqrt{1-(u^{\prime})^{2}}\in W^{1,1}_{\rm loc}(\mathbb{R})$ and $(u^{\prime}/\sqrt{1-(u^{\prime})^{2}})^{\prime}=-\rho\in{\mathcal{C}}^{1}(\mathbb{R})$. Therefore, $u\in{\mathcal{C}}^{2}(\mathbb{R})$.
Once we know the regularity of $u$, it is not difficult to verify the Pohožaev identity. For the details, see [18, Proof of Proposition 2.6]. ∎
3. Concentration-compactness and a profile decomposition
In this section, we establish a profile decomposition for bounded sequences in $X$, which is useful to prove the existence of minimizers of $I$ on ${\mathcal{M}}_{i}$. Recall that in this section, the same notation to Section˜2 is used: $p_{0}$, $X$ and $\Psi_{0}$. For the profile decomposition, several lemmas are necessary. Firstly, a slight extension of [18, Lemma 2.27] is needed:
Lemma 3.1.
Let $Q:=[0,1]^{N}$, $(u_{n})_{n\in\mathbb{N}}\subset X$ be bounded in $X$ and $\zeta\in{\mathcal{C}}(\mathbb{R};[0,\infty))$ satisfy
(3.1) $\limsup_{s\to 0}\frac{\zeta(s)}{|s|^{p_{0}}}=0.$
Then for every $q\in(p_{0},p^{*}_{0})$ and $\varepsilon\in(0,1)$ where $1/p_{0}^{*}=(1/p_{0}-1/N)_{+}$, there exists $C=C(L,\zeta,q,\varepsilon)>0$, where $L:=\sup_{n\in\mathbb{N}}\|u_{n}\|_{X}$, such that
$\int_{\mathbb{R}^{N}}\left|\zeta(u_{n})\right|dx\leq\varepsilon+C\left(\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}\left|u_{n}\right|^{q}dx\right)^{(q-p_{0})/q}.$
In particular, if
$\lim_{n\to\infty}\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}|u_{n}|^{q}\,dx=0,$
then
$\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx\to 0.$
Proof.
Let $q\in(p_{0},p_{0}^{*})$ and $\varepsilon\in(0,1)$. From (2.1) and (2.2), $W^{1,p_{0}}(\mathbb{R}^{N})\subset L^{q}(\mathbb{R}^{N})$ holds. Since $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X$, it is bounded in $L^{\infty}(\mathbb{R}^{N})$. By this bound in $L^{\infty}(\mathbb{R}^{N})$ and (3.1), there exists $C=C(L,\zeta,q,\varepsilon)>0$ such that
$\zeta(u_{n}(x))\leq\varepsilon|u_{n}(x)|^{p_{0}}+C|u_{n}(x)|^{q}\quad\text{for all $x\in\mathbb{R}^{N}$ and $n\in\mathbb{N}$},$
which leads to
(3.2) $\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx\leq\varepsilon\|u_{n}\|_{p_{0}}^{p_{0}}+C\|u_{n}\|_{q}^{q}\quad\text{for all $n\in\mathbb{N}$}.$
Recall $L=\sup_{n\in\mathbb{N}}\|u_{n}\|_{X}<\infty$ and notice also that $W^{1,p_{0}}(Q)\subset L^{q}(Q)$, $X\subset W^{1,p_{0}}(\mathbb{R}^{N})\cap{\mathcal{C}}_{0}(\mathbb{R}^{N})$ and
$\|u_{n}\|_{q}^{q}$ $=\sum_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{q}$
$\leq\left(\sup_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{q-p_{0}}\right)\sum_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{p_{0}}$
$\leq\left(\sup_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{q-p_{0}}\right)\sum_{y\in\mathbb{Z}^{N}}C_{N,p_{0},q}\|u_{n}\|_{W^{1,p_{0}}(y+Q)}^{p_{0}}$
$=C_{N,p_{0},q}\|u_{n}\|_{W^{1,p_{0}}(\mathbb{R}^{N})}^{p_{0}}\sup_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{q-p_{0}}$
$\leq C_{N,p_{0},q}C_{L,N,p_{0}}\sup_{y\in\mathbb{Z}^{N}}\|u_{n}\|_{L^{q}(y+Q)}^{q-p_{0}}.$
Thus (3.2) implies
$\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx\leq\left(\sup_{n\geq 1}\|u_{n}\|_{p_{0}}^{p_{0}}\right)\varepsilon+C\sup_{y\in\mathbb{Z}^{N}}\left(\int_{y+Q}\left|u_{n}\right|^{q}dx\right)^{(q-p_{0})/q}.$
Thus we complete the proof. ∎
Lemma 3.2.
Suppose that $\zeta:\mathbb{R}\rightarrow[0,\infty)$ is of class ${\mathcal{C}}^{1}$ such that $|\zeta^{\prime}(s)|\lesssim|s|^{p_{0}-1}$ for $|s|\leq 1$ and $\zeta(0)=0$. Let $(u_{n})\subset X$ be a bounded sequence such that $u_{n}(x)\to u_{0}(x)$ a.e. $x\in\mathbb{R}^{N}$. Then
$\lim_{n\to\infty}\left(\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx-\int_{\mathbb{R}^{N}}\zeta(u_{n}-u_{0})\,dx\right)=\int_{\mathbb{R}^{N}}\zeta(u_{0})\,dx.$
Proof.
Observe that
$\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx-\int_{\mathbb{R}^{N}}\zeta(u_{n}-u_{0})\,dx$ $=-\int_{\mathbb{R}^{N}}\int_{0}^{1}\frac{d}{ds}\zeta(u_{n}-su_{0})\,ds\,dx=\int_{0}^{1}\int_{\mathbb{R}^{N}}\zeta^{\prime}(u_{n}-su_{0})u_{0}\,dx\,ds$
$\to\int_{0}^{1}\int_{\mathbb{R}^{N}}\zeta^{\prime}(u_{0}-su_{0})u_{0}\,dx\,ds=-\int_{\mathbb{R}^{N}}\int_{0}^{1}\frac{d}{ds}\zeta(u_{0}-su_{0})\,ds\,dx$
$=\int_{\mathbb{R}^{N}}\zeta(u_{0})\,dx-\int_{\mathbb{R}^{N}}\zeta(u_{0}-u_{0})\,dx=\int_{\mathbb{R}^{N}}\zeta(u_{0})\,dx,$
where to pass to the limit we used Vitali’s convergence theorem, $X\subset L^{\infty}(\mathbb{R}^{N})$, the boundedness of $(u_{n})_{n\in\mathbb{N}}$ in $X$ and the estimate $\left|\zeta^{\prime}(u_{n}-su_{0})\right|\lesssim|u_{n}-su_{0}|^{p_{0}-1}$ a.e. $\mathbb{R}^{N}$. ∎
Finally we prove the profile decomposition.
Proposition 3.3.
Let $(u_{n})_{n\in\mathbb{N}}\subset X$ be bounded in $X$. Then there are $K\in\{0,1,2,\ldots\}\cup\{\infty\}$, $(y_{n}^{j})_{n\in\mathbb{N}}\subset\mathbb{R}^{N}$, $\widetilde{u}_{j}\in X$ for $0\leq j<K+1$ such that for any $a\in(0,\infty)$ in the positive mass case,
$u_{n}(\cdot+y_{n}^{j})$ $\rightharpoonup\widetilde{u}_{j}$ $\text{for $0\leq j<K+1$},$
$\widetilde{u}_{j}$ $\neq 0$ $\text{for }1\leq j<K+1,$
$\lim_{n\to\infty}\Psi_{0}(u_{n})$ $\geq\sum_{j=0}^{K}\Psi_{0}(\widetilde{u}_{j})$
and
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{j})\,dx$
for any function $\zeta:\mathbb{R}\rightarrow[0,\infty)$ of class ${\mathcal{C}}^{1}$ such that $|\zeta^{\prime}(s)|\lesssim|s|^{p_{0}-1}$ for $|s|\leq 1$ and
$\limsup_{s\to 0}\frac{\zeta(s)}{|s|^{p_{0}}}=0.$
Proof.
Fix any $q\in(p_{0},p_{0}^{*})$. Since $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X$, up to a subsequence, we may assume $u_{n}\rightharpoonup\widetilde{u}_{0}$ weakly in $X$, $u_{n}\to u_{0}$ a.e. $\mathbb{R}^{N}$ and set
$v^{0}_{n}:=u_{n}-\widetilde{u}_{0},\quad d_{0}:=\limsup_{n\to\infty}\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}|v^{0}_{n}|^{q}\,dx,\quad Q:=[0,1]^{N}.$
If $d_{0}=0$, then we may put $K=0$ and $y_{n}^{0}=0$. The convexity of $\Psi_{0}$ and Sections˜3 and 3 yield
$\liminf_{n\to\infty}\Psi_{0}(u_{n})\geq\Psi_{0}(\widetilde{u}_{0}),\quad\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{0})\,dx+\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(v^{0}_{n})\,dx=\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{0})\,dx.$
Hence the assertions hold.
On the other hand, when $d_{0}>0$, up to a subsequence, we may choose $(y_{n}^{1})_{n\in\mathbb{N}}\subset\mathbb{R}^{N}$ such that
$\int_{y_{n}^{1}+Q}|v^{0}_{n}|^{q}\,dx\to d_{0}>0.$
It is easily seen that $|y_{n}^{1}|\to\infty$ and we may also suppose $v^{0}_{n}(\cdot+y_{n}^{1})\rightharpoonup\widetilde{u}_{1}\not\equiv 0$ weakly in $X$ as well as $v^{0}_{n}(\cdot+y_{n}^{1})\to\widetilde{u}_{1}$ a.e. on $\mathbb{R}^{N}$. Clearly,
$d_{0}=\int_{Q}|\widetilde{u}_{1}|^{q}\,dx.$
We set
$v^{1}_{n}:=v^{0}_{n}-\widetilde{u}_{1}(\cdot-y^{1}_{n})=u_{n}-\sum_{j=0}^{1}\widetilde{u}_{j}(\cdot-y^{j}_{n}),\quad d_{1}:=\limsup_{n\to\infty}\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}|v^{1}_{n}|^{q}\,dx.$
Notice that Section˜3 gives
$\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{0})\,dx+\int_{\mathbb{R}^{N}}\zeta(v^{0}_{n})\,dx+o(1)=\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{0})\,dx+\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{1})\,dx+\int_{\mathbb{R}^{N}}\zeta(v^{1}_{n})\,dx+o(1).$
When $d_{1}=0$, Section˜3 implies
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\sum_{j=0}^{1}\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{j})\,dx.$
For the assertion on $\Psi_{0}$, let $R>0$ be arbitrary and for a domain $\Omega\subset\mathbb{R}^{N}$ write
$\Psi_{0,\Omega}(u):=\begin{dcases}\int_{\Omega}1-\sqrt{1-|\nabla u|^{2}}\,dx&\text{(the zero mass case)},\\ \int_{\Omega}1-\sqrt{1-|\nabla u|^{2}}+au^{2}\,dx&\text{(the positive mass case)}.\end{dcases}$
Since $y^{0}_{n}=0$, $|y_{n}^{1}|\to\infty$ and $u_{n}(\cdot+y^{j}_{n})\rightharpoonup\widetilde{u}_{j}$ weakly in $X$, we obtain
$\liminf_{n\to\infty}\Psi_{0}(u_{n})\geq\liminf_{n\to\infty}\sum_{j=0}^{1}\Psi_{0,B(y^{j}_{n},R)}(u_{n})\geq\sum_{j=0}^{1}\liminf_{n\to\infty}\Psi_{0,B(0,R)}\left(u_{n}(\cdot+y_{n}^{j})\right)\geq\sum_{j=0}^{1}\Psi_{0,B(0,R)}(\widetilde{u}_{j}).$
Letting $R\to\infty$ yields
$\liminf_{n\to\infty}\Psi_{0}(u_{n})\geq\sum_{j=0}^{1}\Psi_{0}(\widetilde{u}_{j})$
and we complete the proof by setting $K=1$.
Now let us consider the case $d_{1}>0$. Since $|y_{n}^{1}|\to\infty$ and $v_{n}^{1}=v_{n}^{0}-\widetilde{u}_{1}(\cdot-y_{n}^{1})$, it is not difficult to see
$d_{1}\leq d_{0}.$
After taking a subsequence if necessary, we may choose $(y^{2}_{n})_{n\in\mathbb{N}}\subset\mathbb{R}^{N}$ so that
$\int_{y^{2}_{n}+Q}|v^{1}_{n}|^{q}\,dx\to d_{1}>0.$
Then $|y^{2}_{n}-y_{n}^{j}|\to\infty$ for $j=0,1$ and $v^{1}_{n}(\cdot+y^{2}_{j})\rightharpoonup\widetilde{u}_{2}\not\equiv 0$ weakly in $X$ as well as $v^{1}_{n}(\cdot+y^{2}_{n})\to\widetilde{u}_{2}$ a.e. $\mathbb{R}^{N}$. Moreover, $\int_{Q}\left|\widetilde{u}_{2}\right|^{q}dx=d_{1}$ holds. We define
$v_{n}^{2}:=v_{n}^{1}-\widetilde{u}_{2}(\cdot-y^{2}_{n}),\quad d_{2}:=\limsup_{n\to\infty}\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}|v_{n}^{2}|^{q}\,dx.$
Section˜3 gives
$\int_{\mathbb{R}^{N}}\zeta(v_{n}^{1})\,dx=\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{2})\,dx+\int_{\mathbb{R}^{N}}\zeta(v_{n}^{2})\,dx+o(1),\quad\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\sum_{j=0}^{2}\zeta(\widetilde{u}_{j})\,dx+\int_{\mathbb{R}^{N}}\zeta(v_{n}^{2})\,dx+o(1).$
If $d_{2}=0$, then we set $K=2$ and may prove the assertions similarly in the case $K=1$.
On the other hand, if $d_{2}>0$, then by $|y_{n}^{1}|\to\infty$, $|y_{n}^{1}-y_{n}^{2}|\to\infty$ and $v^{2}_{n}=v_{n}^{1}-\widetilde{u}_{2}(\cdot-y_{n}^{2})$, $d_{2}\leq d_{1}$ holds. Furthermore, we find $(y^{3}_{n})_{n\in\mathbb{N}}$ and $\widetilde{u}_{3}\not\equiv 0$ such that
$v^{2}_{n}(\cdot+y_{n}^{3})\rightharpoonup\widetilde{u}_{3}\not\equiv 0,\quad d_{2}=\lim_{n\to\infty}\int_{y^{3}_{n}+Q}|v^{2}_{n}|^{q}\,dx=\int_{Q}|\widetilde{u}_{3}|^{q}\,dx,\quad|y_{n}^{3}-y_{n}^{j}|\to\infty\ (j=0,1,2).$
By setting
$v^{3}_{n}:=v^{2}_{n}-\widetilde{u}_{3}(\cdot-y^{3}_{n}),\quad d_{3}:=\limsup_{n\to\infty}\sup_{y\in\mathbb{Z}^{N}}\int_{y+Q}|v^{3}_{n}|^{q}\,dx,$
we repeat the same procedure. Thus, may define $(d_{k})$, $(v^{k}_{n})_{n\in\mathbb{N}}$, $(y^{k}_{n})_{n\in\mathbb{N}}$ and $(\widetilde{u}_{j})_{0\leq j\leq k}$ for each $k\geq 1$. If $d_{k}=0$ holds for some $k\in\mathbb{N}$, then the procedure is terminated and the assertions hold for $K:=k$. Therefore, it suffices to consider the case $d_{k}>0$ for all $k\geq 1$. Remark that for any $k\in\mathbb{N}$,
$\int_{Q}|\widetilde{u}_{k+2}|^{q}\,dx=d_{k+1}\leq d_{k}=\int_{Q}|\widetilde{u}_{k+1}|^{q}\,dx.$
We next claim that
(3.3) $\sum_{j=0}^{\infty}d_{j}<\infty.$
In fact, by the classical Brezis–Lieb lemma (or Section˜3), for every $k\in\mathbb{N}$, we have
$\|u_{n}\|_{q}^{q}=\sum_{j=0}^{k+1}\|\widetilde{u}_{j}\|_{q}^{q}+\|v^{k+1}_{j}\|_{q}^{q}+o(1).$
Therefore, by $q\in(p_{0},p_{0}^{*})$, for each $k\in\mathbb{N}$,
$\sum_{j=0}^{k+1}d_{j}=\sum_{j=0}^{k}\int_{Q}|\widetilde{u}_{j+1}|^{q}\,dx\leq\liminf_{n\to\infty}\left[\|u_{n}\|_{q}^{q}-\|\widetilde{u}_{0}\|_{q}^{q}-\|v^{k+1}_{j}\|_{q}^{q}\right]\leq\sup_{n\geq 1}\|u_{n}\|_{q}^{q}<\infty.$
Since $k\in\mathbb{N}$ is arbitrary, (3.3) holds. In particular, $d_{k}\to 0$ holds as $k\to\infty$.
To apply Section˜3 for $(v^{k}_{n})_{n\in\mathbb{N}}$ uniformly with respect to $k\geq 1$, we prove that for every $k\in\mathbb{N}$,
(3.4) $\limsup_{n\to\infty}\|v^{k}_{n}\|_{X}\leq 2\limsup_{n\to\infty}\|u_{n}\|_{X}.$
For this purpose, let $k\geq 1$. Since
$v^{k}_{n}(x)=u_{n}(x)-\sum_{j=0}^{k}\widetilde{u}_{j}(x-y_{n}^{j}),\quad|j_{1}-j_{2}|\to\infty\ \ (j_{1}\neq j_{2}),$
it follows that
$\left\|\sum_{j=0}^{k}\widetilde{u}_{j}(\cdot-y_{n}^{j})\right\|_{X}\to\sum_{j=0}^{k}\|\widetilde{u}_{j}\|_{X}.$
By a similar argument to $\Psi_{0}$, we get
$\sum_{j=0}^{k}\|\widetilde{u}_{j}\|_{X}\leq\liminf_{n\to\infty}\|u_{n}\|_{X}.$
Thus,
$\limsup_{n\to\infty}\|v^{k}_{n}\|_{X}\leq\limsup_{n\to\infty}\|u_{n}\|_{X}+\limsup_{n\to\infty}\left\|\sum_{j=0}^{k}\widetilde{u}_{j}(\cdot-y^{j}_{n})\right\|_{X}\leq 2\limsup_{n\to\infty}\|u_{n}\|_{X}$
and (3.4) holds.
By Section˜3, (3.4) and
$\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\sum_{j=0}^{k}\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{j})\,dx+\int_{\mathbb{R}^{N}}\zeta(v^{k}_{n})\,dx+o(1),$
for any $\varepsilon>0$ there exist $C_{\varepsilon}>0$ and $\theta=(q-p_{0})/q\in(0,1)$, which does not depend on $k$, such that for each $k\geq 1$,
$\limsup_{n\to\infty}\int_{\mathbb{R}^{N}}\left|\zeta(v^{k}_{n})\right|dx\leq\varepsilon+C_{\varepsilon}d_{k}^{\theta}.$
From $d_{k}\to 0$ as $k\to\infty$, choose $k_{\varepsilon}\in\mathbb{N}$ so large that $C_{\varepsilon}d_{k}^{\theta}\leq\varepsilon$ for all $k\geq k_{\varepsilon}$. Thus, if $k\geq k_{\varepsilon}$, then
$\limsup_{n\to\infty}\left|\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx-\sum_{j=0}^{k}\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{j})\,dx\right|\leq 2\varepsilon.$
Since we may assume that the limit $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx$ exists, the above inequality implies
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}\zeta(u_{n})\,dx=\sum_{j=0}^{\infty}\int_{\mathbb{R}^{N}}\zeta(\widetilde{u}_{j})\,dx.$
For the assertion on $\Psi_{0}$, for any $k\in\mathbb{N}$ and $R>0$, we may prove
$\liminf_{n\to\infty}\Psi_{0}(u_{n})\geq\sum_{j=0}^{k}\Psi_{0,B(0,R)}(\widetilde{u}_{j}).$
By letting $R\to\infty$ and then $k\to\infty$, the desired assertion holds. ∎
4. The zero mass case
In this section, the zero mass case is considered and Theorems˜1.2 and 1.4 are proved. From now on, $N\geq 3$ and (F1)–(F6) are always supposed. We first investigate properties of $M_{0}$ and ${\mathcal{M}}_{0}$ defined in (1.6) and (1.7), and then prove Theorems˜1.2 and 1.4.
4.1. The projection into the Pohožaev set
For $u\in D(\Psi)\setminus\{0\}$ and $\theta>0$, we consider the scaling
$u_{\theta}:=\theta u\left(\frac{\cdot}{\theta}\right).$
It is immediate to see
$\|\nabla u_{\theta}\|_{q}^{q}=\theta^{N}\|\nabla u\|_{q}^{q}\quad\mbox{ for }q\geq 1,\quad\|\nabla u_{\theta}\|_{\infty}=\|\nabla u\|_{\infty}.$
From the above observation we see that $u_{\theta}\to 0$ in $X_{0}$ as $\theta\to 0$. By a direct calculation we obtain
(4.1) $I(u_{\theta})$ $=\theta^{N}\left(\Psi(u)-\Phi(\theta u)\right),$
$\frac{d}{d\theta}I(u_{\theta})$ $=N\theta^{N-1}\left(\Psi(u)-\Phi(\theta u)-\frac{1}{N}\Phi^{\prime}(\theta u)(\theta u)\right)=N\theta^{N-1}\left(\Psi(u)-\int_{\mathbb{R}^{N}}H(\theta u)\,dx\right).$
To study properties of the Pohožaev set ${\mathcal{M}}_{0}$, we first prove
Lemma 4.1.
For each $\alpha\in(0,\infty)$ and $u\in X_{0}\setminus\{0\}$, there exists a unique $\Theta=\Theta(\alpha,u)\in(0,\infty)$ such that
$\alpha=\int_{\mathbb{R}^{N}}H(\Theta u)\,dx.$
Moreover, $\Theta\in{\mathcal{C}}^{1}((0,\infty)\times X_{0}\setminus\{0\};(0,\infty))$.
Proof.
Define
$J(\alpha,u,\theta):=\alpha-\int_{\mathbb{R}^{N}}H(\theta u)\,dx:(0,\infty)\times X_{0}\setminus\{0\}\times(0,\infty)\to\mathbb{R}.$
In view of $X_{0}\subset L^{\infty}(\mathbb{R}^{N})$ and (F1)–(F2), $J$ is of class ${\mathcal{C}}^{1}$. For any fixed $u\in X_{0}\setminus\{0\}$, it follows from (F3) that
(4.2) $\frac{d}{d\theta}\int_{\mathbb{R}^{N}}H(\theta u)\,dx=\theta^{-1}\int_{\mathbb{R}^{N}}H^{\prime}(\theta u)\theta u\,dx>0.$
Furthermore, by (F2)–(F4) and $u\in X_{0}\setminus\{0\}$,
$\lim_{\theta\to 0}\int_{\mathbb{R}^{N}}H(\theta u)\,dx=0,\quad\lim_{\theta\to\infty}\int_{\mathbb{R}^{N}}H(\theta u)\,dx=\infty.$
Therefore, for each $\alpha\in(0,\infty)$ and $u\in X_{0}\setminus\{0\}$, there exists a unique $\Theta=\Theta(\alpha,u)\in(0,\infty)$ such that $J(\alpha,u,\Theta)=0$. Since
$\frac{\partial J}{\partial\theta}(\alpha,u,\Theta)=-\Theta^{-1}\int_{\mathbb{R}^{N}}H^{\prime}(\Theta u)\Theta u\,dx<0,$
the regularity of $\Theta$ follows from the implicit function theorem. ∎
By Sections˜2 and 4.1, we may construct a continuous projection from $D(\Psi)\setminus\{0\}$ onto ${\mathcal{M}}_{0}$:
Lemma 4.2.
For every $u\in D(\Psi)\setminus\{0\}$ there is unique $\theta=\theta(u)>0$ such that $u_{\theta}\in{\mathcal{M}}_{0}$. Moreover the map
$m(u):=u_{\theta(u)}:D(\Psi)\setminus\{0\}\to{\mathcal{M}}_{0}$
is continuous, where the topology on $D(\Psi)\setminus\{0\}$ is induced from $X_{0}$.
Proof.
Recall (1.6) and (1.7). Since
$M_{0}(u_{\theta})=\theta^{N}\left(\Psi(u)-\int_{\mathbb{R}^{N}}H(\theta u)\,dx\right),$
the desired map $m$ is given by setting $\theta(u)\coloneq\Theta(\Psi(u),u)$ where $\Theta$ appears in Subsection˜4.1. Since $\Psi:D(\Psi)\setminus\{0\}\to\mathbb{R}$ is continuous by Section˜2, $\theta(\cdot)$ is continuous, and hence so is $m$. ∎
Remark 4.3.
Let $u\in D(\Psi)\setminus\{0\}$. From (4.1) and (4.2), $\theta=\theta(u)$ is the only maximum of the function $0,\infty)\ni\theta\mapsto I(u_{\theta})$ with $I(u_{\theta(u)})>0$.
4.2. The ground state level
In this subsection we are going to show $c_{0}=c_{MP}>0$ where $c_{0,MP}$ is the mountain pass level
(4.3) $c_{0,MP}:=\inf_{\gamma\in\Gamma}\sup_{\tau\in[0,1]}I(\gamma(\tau)),\quad\Gamma:=\left\{\gamma\in{\mathcal{C}}([0,1];X_{0})\ :\ \gamma(0)=0,\ I(\gamma(1))<0\right\}.$
Remark 4.4.
Fix any $\varphi\in{\mathcal{C}}^{\infty}_{0}(\mathbb{R}^{N})$ so that $\|\nabla\varphi\|_{\infty}<1$. Then for each $\theta>0$, $\varphi_{\theta}=\theta\varphi(\cdot/\theta)\in D(\Psi)$ holds. Moreover, as in Subsection˜4.1, we may prove $\|\varphi_{\theta}\|_{X_{0}}\to 0$ as $\theta\to 0$ and $I(\varphi_{\theta})\to$-$\infty$ as $\theta\to\infty$. Therefore, for a sufficiently large $T>0$, $\gamma_{0}(\theta):=\varphi_{T\theta}\in\Gamma$ and $c_{0,MP}<\infty$ holds. Moreover, it is clear that
$c_{0,MP}=\inf_{\gamma\in\widetilde{\Gamma}}\sup_{\tau\in[0,1]}I(\gamma(\tau)),\quad\widetilde{\Gamma}:=\left\{\gamma\in{\mathcal{C}}([0,1];X_{0})\ :\ \gamma([0,1])\subset D(\Psi),\ \gamma(0)=0,\ I(\gamma(1))<0\right\}.$
We also remark that for any $\gamma\in\widetilde{\Gamma}$, the function $0,1]\ni t\mapsto I(\gamma(t))\in\mathbb{R}$ is continuous by virtue of Section˜2.
Lemma 4.5.
There holds $c_{0}=c_{0,MP}>0$.
Proof.
We first prove $c_{0}\geq c_{0,MP}$. Fix any $u\in{\mathcal{M}}_{0}$. Then we know that $\theta(u)=1$. Consider a path $\gamma_{u}(\tau):=u_{T_{u}\tau}$, where $T_{u}\gg 1$ and $\gamma_{u}(0):=\lim_{\tau\to 0^{+}}u_{T_{u}\tau}=0$. Note that $I(u_{T_{u}})<0$ for sufficiently large $T_{u}$, hence $\gamma_{u}\in\Gamma$. Since Subsection˜4.1 leads to $c_{0,MP}\leq\max_{\tau\in[0,1]}I(\gamma_{u}(\tau))=I(u)$ and $u\in{\mathcal{M}}_{0}$ is arbitrary,
$c_{0,MP}\leq\inf_{u\in{\mathcal{M}}_{0}}I(u)=c_{0}.$
To show $c_{0,MP}>0$ we recall that $X_{0}\subset L^{\infty}(\mathbb{R}^{N})$ and $|F(s)|\lesssim|s|^{2^{*}}$ for $|s|\leq 1$ thanks to (F2). By (2.4), if $\|u\|_{X_{0}}$ is sufficiently small, then
$I(u)$ $\geq\sum_{j=1}^{\infty}b_{j}\|\nabla u\|_{2j}^{2j}-C\|u\|_{2^{*}}^{2^{*}}$
$\geq b_{1}\|\nabla u\|_{2}^{2}+b_{j_{0}}\|\nabla u\|_{2j_{0}}^{2j_{0}}-C\|\nabla u\|_{2}^{2^{*}}$
$=\left(b_{1}-C\|\nabla u\|_{2}^{2^{*}-2}\right)\|\nabla u\|_{2}^{2}+b_{j_{0}}\|\nabla u\|_{2j_{0}}^{2j_{0}}$
$\gtrsim\|u\|^{2j_{0}}_{X_{0}}.$
Hence there is a small radius $\rho>0$ such that $\inf_{\|u\|_{X_{0}}=\rho_{0}}I(u)>0$ and $I(u)\geq 0$ provided $\|u\|_{X_{0}}\leq\rho_{0}$. Thus, $c_{0,MP}>0$.
We next prove $c_{0}\leq c_{0,MP}$. By Subsection˜4.2, as in [23], it suffices to show $\gamma([0,1])\cap{\mathcal{M}}_{0}\neq\emptyset$ for any $\gamma\in\widetilde{\Gamma}$. Let $\gamma\in\widetilde{\Gamma}$ and $t_{0}:=\max\{t\in[0,1]\ :\ \gamma(t)=0\}\in[0,1)$. For $M_{0}$ in (1.7), as in the above, for sufficiently small $\delta>0$, we observe that
$M_{0}(\gamma(t))>0\quad\text{for all $t\in(t_{0},t_{0}+\delta)$}.$
On the other hand, since (F5) implies $f(s)s\geq 0$ for any $s\in\mathbb{R}$, it follows that $\Phi^{\prime}(u)u\geq 0$ for every $u\in X_{0}$ and
$M_{0}(\gamma(1))=I(\gamma(1))-\frac{1}{N}\Phi^{\prime}(\gamma(1))\gamma(1)\leq I(\gamma(1))<0.$
Since the function $0,1]\ni t\mapsto M_{0}(\gamma(t))\in\mathbb{R}$ is continuous by Section˜2, there exists $t_{1}\in(t_{0},1)$ such that $M_{0}(\gamma(t_{1}))=0$, which yields $\gamma(t_{1})\in{\mathcal{M}}_{0}$ and $\gamma([0,1])\cap{\mathcal{M}}_{0}\neq\emptyset$. Thus we complete the proof. ∎
4.3. Minimizers on the Pohožaev set are critical points
The aim of this subsection is to prove that every minimizer corresponding to $c_{0}$ is a critical point of $I$, and hence Section˜2 implies that it is a classical solution of (1.1).
Lemma 4.6.
Let $u_{0}\in{\mathcal{M}}_{0}$ satisfy $I(u_{0})=c_{0}$. Then $u_{0}$ is a critical point of $I$.
Proof.
Suppose that $u_{0}\in{\mathcal{M}}$ satisfies $I(u_{0})=c_{0}>0$. By Section˜2, there is $r_{0}>0$ such that
$\frac{1}{2}\Psi(u_{0})\leq\Psi(u)\leq 2\Psi(u_{0})$
for $u\in\overline{B_{X_{0}}(u_{0},r_{0})}\cap D(\Psi)$. We recall that
$\frac{d}{d\theta}I(u_{\theta})=N\theta^{N-1}\left(\Psi(u)-\int_{\mathbb{R}^{N}}H(\theta u)\,dx\right)$
and therefore we have the following formula
$\frac{d^{2}}{d\theta^{2}}I(u_{\theta})=N(N-1)\theta^{N-2}\left(\Psi(u)-\int_{\mathbb{R}^{N}}H(\theta u)\,dx\right)-N\theta^{N-2}\int_{\mathbb{R}^{N}}H^{\prime}(\theta u)(\theta u)\,dx.$
Hence
$L_{1}:=\sup\left\{\left|\frac{d^{2}}{d\theta^{2}}I(u_{\theta})\right|\ :\ \frac{1}{2}\leq\theta\leq 2,\ u\in\overline{B_{X}(u_{0},r_{0})}\cap D(\Psi)\right\}<\infty.$
From $dI(u_{\theta})/d\theta|_{\theta=\theta(u)}=0$, $\theta(u_{0})=1$ and the continuity of $\theta(\cdot)$, it follows that for $u\in\overline{B_{X_{0}}(u_{0},r_{0})}\cap D(\Psi)$,
(4.4) $\frac{1}{2}\leq\theta(u)\leq 2,\quad\left|I(u_{\theta(u)})-I(u)\right|\leq\left|\int_{\theta(u)}^{1}\frac{d}{d\theta}I(u_{\theta})\,d\theta\right|\leq\frac{L_{1}}{2}|1-\theta(u)|^{2}.$
To prove that $u_{0}$ is a critical point of $I$, let $v\in D(\Psi)$ be any point: otherwise, (1.5) holds trivially. For $t\in(0,1)$, we introduce
$v_{t}:=(1-t)u_{0}+tv\in D(\Psi).$
The convexity of $\Psi$ implies that for any $t\in(0,1]$,
(4.5) $\Psi(v_{t})-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v_{t}-u_{0})\leq t\left\{\Psi(v)-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v-u_{0})\right\}.$
Let $(t_{n})\subset(0,1)$ satisfy $t_{n}\to 0$. We first consider the case $\Psi(u_{0})\leq\Psi(v_{t_{n}})$ for all $n$. Then it is clear that $v_{t_{n}}\to u_{0}$ in $X_{0}$ and from the convexity of $\Psi$,
(4.6) $0\leq\Psi(v_{t_{n}})-\Psi(u_{0})\leq(1-t_{n})\Psi(u_{0})+t_{n}\Psi(v)-\Psi(u_{0})=t_{n}\left(\Psi(v)-\Psi(u_{0})\right).$
To simplify notations we put
$\theta_{n}:=\theta(v_{t_{n}}),\quad w_{n}:=(v_{t_{n}})_{\theta_{n}}=\theta_{n}v_{t_{n}}\left(\frac{\cdot}{\theta_{n}}\right)\in{\mathcal{M}}_{0}.$
From $\|v_{t_{n}}-u_{0}\|_{X_{0}}=O(t_{n})$ and (4.6) with $\Theta\in{\mathcal{C}}^{1}((0,\infty)\times X_{0}\setminus\{0\})$ and $\Theta(\Psi(u),u)=\theta(u)$ for $u\in D(\Psi)\setminus\{0\}$, we get
(4.7) $|\theta_{n}-1|=|\Theta(\Psi(v_{t_{n}}),v_{t_{n}})-\Theta(\Psi(u_{0}),u_{0})|=O(t_{n}).$
Using the fact that $\Phi\in{\mathcal{C}}^{1}(X;\mathbb{R})$ and (4.5), we infer that
(4.8) $I(v_{t_{n}})-I(u_{0})$ $=\Psi(v_{t_{n}})-\Psi(u_{0})-(\Phi(v_{t_{n}})-\Phi(u_{0}))$
$=\Psi(v_{t_{n}})-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v_{t_{n}}-u_{0})+o(t_{n})$
$\leq t_{n}\left\{\Psi(v)-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v-u_{0})\right\}+o(t_{n}).$
By (4.4) and (4.7) with $v_{t_{n}}\in\overline{B_{X_{0}}(u_{0},r_{0})}\cap D(\Psi)$,
$I(w_{n})-I(v_{t_{n}})\leq\frac{L_{1}}{2}|\theta_{n}-1|^{2}=O(t_{n}^{2})$
and recalling that $u_{0}$ minimizes $I$ on ${\mathcal{M}}_{0}$, we see from (4.8) that
$0\leq I(w_{n})-I(u_{0})\leq I(v_{t_{n}})-I(u_{0})+O(t_{n}^{2})\leq t_{n}\left\{\Psi(v)-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v-u_{0})\right\}+o(t_{n}).$
Dividing the above inequality by $t_{n}$ and sending $t_{n}\to 0$ imply (1.5) under $\Psi(u_{0})\leq\Psi(v_{t_{n}})$ for each $n$.
If there exists a subsequence $(n_{k})_{k\in\mathbb{N}}$ such that $\Psi(u_{0})\leq\Psi(v_{t_{n_{k}}})$ for all $k$, then the previous argument is still valid. Therefore, we may assume $\Psi(u_{0})>\Psi(v_{t_{n}})$ for all $n$. We will use the same notation as in the previous case. If
$\limsup_{n\to\infty}\frac{|\theta_{n}-1|}{t_{n}}<\infty,$
then (4.7) holds in this case and the previous argument can be exploited to obtain
$\Psi(v)-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v-u)\geq 0.$
Consequently, we may assume that
(4.9) $\lim_{n\to\infty}\frac{|\theta_{n}-1|}{t_{n}}=\infty.$
In what follows, we shall show that (4.9) does not occur, which implies that the desired inequality is satisfied by any $v\in D(\Psi)$ and this completes the proof. For this purpose, we next show $\theta_{n}\in(0,1)$ for all sufficiently large $n$. Indeed, suppose that there exists $(\theta_{n_{k}})$ such that $n_{k}\to\infty$ and $\theta_{n_{k}}\geq 1$. By $\|v_{t_{n_{k}}}-u_{0}\|_{X_{0}}=ct_{n_{k}}$ and
$0=\Psi(v_{t_{n_{k}}})-\int_{\mathbb{R}^{N}}H(\theta_{n_{k}}v_{t_{n_{k}}})\,dx,$
it follows from the facts $u_{0}\in{\mathcal{M}}_{0}$ and $\Psi(u_{0})>\Psi(v_{t_{n}})$ in view of the assumption that for some $C_{0}>0$, which is independent of $k$,
$\int_{\mathbb{R}^{N}}H(u_{0})\,dx=\Psi(u_{0})>\Psi(v_{t_{n_{k}}})=\int_{\mathbb{R}^{N}}H(\theta_{n_{k}}v_{t_{n_{k}}})\,dx\geq\int_{\mathbb{R}^{N}}H(\theta_{n_{k}}u_{0})\,dx-C_{0}t_{n_{k}}.$
Thus, we may find $\tau_{k}\in(0,1)$ such that
(4.10) $C_{0}t_{n_{k}}\geq\int_{\mathbb{R}^{N}}H(\theta_{n_{k}}u_{0})-H(u_{0})\,dx=\int_{\mathbb{R}^{N}}H^{\prime}\left(\left(1+\tau_{k}(\theta_{n_{k}}-1)\right)u_{0}\right)u_{0}\,dx\left(\theta_{n_{k}}-1\right).$
However, since $\theta_{n}=\theta(v_{t_{n}})\to 1$ and (F3) lead to
$\lim_{k\to\infty}\int_{\mathbb{R}^{N}}H^{\prime}\left(\left(1+\tau_{k}(\theta_{n_{k}}-1)\right)u_{0}\right)u_{0}\,dx=\int_{\mathbb{R}^{N}}H^{\prime}(u_{0})u_{0}\,dx>0,$
(4.10) contradicts (4.9). Therefore, $\theta_{n}\in(0,1)$ holds for all sufficiently large $n$.
Since $\theta_{n}<1$ and $I(u)=\frac{1}{N}\int_{\mathbb{R}^{N}}f(u)u\,dx$ for $u\in{\mathcal{M}}_{0}$, (F5) and the fact $w_{n}\in{\mathcal{M}}_{0}$ give
$I(u_{0})\leq I(w_{n})$ $=\frac{\theta_{n}^{N}}{N}\int_{\mathbb{R}^{N}}f(\theta_{n}v_{t_{n}})\theta_{n}v_{t_{n}}\,dx$
(4.11) $\leq\frac{\theta_{n}^{N}}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx$
$=\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx-(1-\theta_{n}^{N})\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx.$
Observe that
$\lim_{n\to\infty}\frac{1-\theta_{n}^{N}}{t_{n}}\geq\lim_{n\to\infty}\frac{1-\theta_{n}}{t_{n}}=\infty$
and that
$\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\to\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx=NI(u_{0})=Nc_{0}>0.$
Moreover by $\|v_{t_{n}}-u_{0}\|_{X_{0}}=t_{n}\|v-u_{0}\|_{X_{0}}=O(t_{n})$ and the fact that $X_{0}\ni u\mapsto\int_{\mathbb{R}^{N}}f(u)u\,dx\in\mathbb{R}$ is of class ${\mathcal{C}}^{1}$, for some $C_{1}>0$,
$\left|NI(u_{0})-\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\right|=\left|\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx-\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\right|\leq C_{1}t_{n}.$
However this contradicts (4.11). Indeed, from (4.11) we obtain
$I(u_{0})-\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\leq-(1-\theta_{n}^{N})\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx$
or equivalently
$(1-\theta_{n}^{N})\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\leq\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx-I(u_{0})\leq\frac{C_{1}}{N}t_{n}.$
This yields the following contradiction:
$0<\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx=\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\leq C_{1}\limsup_{n\to\infty}\frac{t_{n}}{1-\theta_{n}^{N}}=0.$
Thus (4.9) does not occur and $u_{0}$ is a critical point of $I$. ∎
4.4. Existence of a minimizer and other properties of minimizers
To prove Theorem˜1.2, from Sections˜2, 4.2 and 4.3, it suffices to prove that there exists a minimizer, any minimizer has a constant sign and there exists a radial minimizer corresponding to $c_{0}$. We will show the existence of a minimizer by a direct minimization technique on ${\mathcal{M}}_{0}$ and the profile decomposition.
Proof of Theorem˜1.2.
Let $(u_{n})_{n\in\mathbb{N}}\subset{\mathcal{M}}_{0}$ be a minimizing sequence, namely $I(u_{n})\to c_{0}$. Note that for $u\in{\mathcal{M}}_{0}$ we have
$I(u)=\frac{1}{N}\int_{\mathbb{R}^{N}}f(u)u\,dx.$
Thus, $(\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx)_{n\in\mathbb{N}}$ is bounded. From (F6), $(\int_{\mathbb{R}^{N}}F(u_{n})\,dx)_{n\in\mathbb{N}}$ is bounded and by $I(u_{n})\to c_{0}$, $(\Psi(u_{n}))_{n\in\mathbb{N}}$ is also bounded. Hence $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X_{0}$ via (2.4). From Section˜3 with $p_{0}=2^{*}$, there exist $K\in\{0,1,2,\dots\}\cup\{\infty\}$, $(y_{n}^{j})_{n}$ and $\widetilde{u}_{j}$ such that
(4.12) $u_{n}(\cdot+y_{n}^{j})\rightharpoonup\widetilde{u}_{j},\quad u_{n}(x+y_{n}^{j})\to\widetilde{u}_{j}(x)\quad\text{a.e. $\mathbb{R}^{N}$},$
$\lim_{n\to\infty}\Psi(u_{n})\geq\sum_{j=0}^{K}\Psi(\widetilde{u}_{j}),$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}F(u_{n})\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}F(\widetilde{u}_{j})\,dx,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j})\widetilde{u}_{j}\,dx.$
If $K=0$ and $\widetilde{u}_{0}=0$, then
$I(u_{n})=\frac{1}{N}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx\to 0$
and $c_{0}=0$, which is a contradiction. Hence either $K\geq 1$ or $K=0$ and $\widetilde{u}_{0}\not\equiv 0$.
Observe that ${\mathcal{M}}_{0}\subset D(\Psi)\setminus\{0\}$ due to $M_{0}(u)=\infty$ for $u\not\in D(\Psi)$, hence $u_{n}\in D(\Psi)\setminus\{0\}$. Since $D(\Psi)$ is convex and closed, it is also weakly closed thanks to Mazur’s theorem, which gives $\widetilde{u}_{j}\in D(\Psi)$.
To treat the cases $K\geq 1$ and $K=0$ with $\widetilde{u}_{0}\not\equiv 0$ simultaneously, set $J:=\left\{j\ :\ \widetilde{u}_{j}\not\equiv 0\right\}$. Remark that $J\neq\emptyset$ by the above observation. For $j\in J$, let $\theta_{j}>0$ be the unique number so that $(\widetilde{u}_{j})_{\theta_{j}}\in{\mathcal{M}}_{0}$. Suppose that for all $j\in J$ we have $\theta_{j}>1$. By (4.2) and (4.12),
(4.13) $0<\sum_{j\in J}\Psi(\widetilde{u}_{j})\leq\lim_{n\to\infty}\Psi(u_{n})=\lim_{n\to\infty}\int_{\mathbb{R}^{N}}H(u_{n})\,dx$ $=\sum_{j\in J}\int_{\mathbb{R}^{N}}H(\widetilde{u}_{j})\,dx$
$<\sum_{j\in J}\int_{\mathbb{R}^{N}}H(\theta_{j}\widetilde{u}_{j})\,dx=\sum_{j\in J}\Psi(\widetilde{u}_{j}),$
which is a contradiction. Hence there exists $j_{0}\in J$ such that $\theta_{j_{0}}\leq 1$. Then, Fatou’s lemma and (F5) yield
(4.14) $c_{0}=\lim_{n\to\infty}I(u_{n})$ $=\lim_{n\to\infty}\frac{1}{N}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\lim_{n\to\infty}\frac{1}{N}\int_{\mathbb{R}^{N}}f(u_{n}(\cdot+y_{n}^{j_{0}}))u_{n}(\cdot+y_{n}^{j_{0}})\,dx$
$\geq\frac{1}{N}\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j_{0}})\widetilde{u}_{j_{0}}\,dx\geq\frac{1}{N}\theta_{j_{0}}^{N}\int_{\mathbb{R}^{N}}f(\theta_{j_{0}}\widetilde{u}_{j_{0}})\theta_{j_{0}}\widetilde{u}_{j_{0}}\,dx=I(m(\widetilde{u}_{j_{0}}))\geq c_{0}.$
Thus $\theta_{j_{0}}=1$, $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{0}$, $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j_{0}})\widetilde{u}_{j_{0}}\,dx$ and $I(\widetilde{u}_{j_{0}})=c_{0}$. Moreover, since $f(s)s>0$ for $s\neq 0$ in view of (F3) and (F5), by (4.12), $\widetilde{u}_{j}=0$ for each $j\neq j_{0}$, $J=\{j_{0}\}$ and $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}F(u_{n})\,dx=\int_{\mathbb{R}^{N}}F(\widetilde{u}_{j_{0}})\,dx$. Furthermore, by $I(u_{n})\to c_{0}=I(\widetilde{u}_{j_{0}})$, we have $\Psi(u_{n})\to\Psi(\widetilde{u}_{j_{0}})$, and Section˜2 yields $\|u_{n}(\cdot+y_{n}^{j_{0}})-\widetilde{u}_{j_{0}}\|_{X_{0}}\to 0$. Thanks to Subsection˜4.3 we know that $\widetilde{u}_{j_{0}}$ is a critical point of $I$ and Section˜2 implies that $\widetilde{u}_{j_{0}}$ is a classical solution of (1.1).
We next prove that all minimizers have a constant sign. To this end, let $u\in{\mathcal{M}}_{0}$ be any minimizer corresponding to $c_{0}$. We argue by contradiction and suppose $u_{\pm}\not\equiv 0$ where $u_{+}:=\max\{u,0\}$ and $u_{-}:=\min\{u,0\}$. From $\|\nabla u\|_{\infty}<1$ due to Sections˜2 and 4.3, it follows that $u_{+},u_{-}\in D(\Psi)\setminus\{0\}$. According to Subsection˜4.1 there exist $\theta_{+}$ and $\theta_{-}$ such that $(u_{+})_{\theta_{+}},(u_{-})_{\theta_{-}}\in{\mathcal{M}}_{0}$. By
$\Psi(u_{+})+\Psi(u_{-})=\Psi(u)=\int_{\mathbb{R}^{N}}H(u)\,dx=\int_{\mathbb{R}^{N}}H(u_{+})\,dx+\int_{\mathbb{R}^{N}}H(u_{-})\,dx,$
arguing as in (4.13), we get either $\theta_{+}\leq 1$ or $\theta_{-}\leq 1$. Let $\theta_{+}\leq 1$. By noting $f(s)s\geq 0$ and replacing $u_{n}$ and $\widetilde{u}_{j_{0}}$ to $u$ and $u_{+}$ in (4.14), we deduce that $\theta_{+}=1$, $u_{+}\in{\mathcal{M}}_{0}$, $I(u_{+})=c_{0}$ and $u_{-}\equiv 0$, which is a contradiction. Similarly, a contradiction occurs when $\theta_{-}\leq 1$ holds. Therefore, either $u_{+}\equiv 0$ or $u_{-}\equiv 0$. Since $u$ is a classical solution of (1.1) with either $u\geq 0$ or $u\leq 0$, the strong maximum principle gives either $u>0$ or $u<0$ and thus $u$ has a constant sign.
Finally, we prove the existence of a radial minmizer corresponding to $c_{0}$. Let $u\in{\mathcal{M}}_{0}$ be arbitrary minimizer of $c_{0}$. From the above, either $u>0$ or $u<0$ holds. Set $v:=u^{*}$ when $u>0$ where $u^{*}$ is the symmetric decreasing rearrangement of $u$, and $v:=-u^{*}$ when $u<0$. From [33, Lemma 1] and (2.4) (or [12, Lemma 4.3]) it follows that
$\Psi(u)=\sum_{j\in\mathbb{N}}b_{j}\|\nabla u\|_{2j}^{2j}\geq\sum_{j\in\mathbb{N}}b_{j}\|\nabla u_{0}^{*}\|_{2j}^{2j}=\Psi(v).$
On the other hand, since $u$ has a constant sign and $f\in{\mathcal{C}}^{1}$, we also have
$\int_{\mathbb{R}^{N}}F(u)\,dx=\int_{\mathbb{R}^{N}}F(v)\,dx,\quad\int_{\mathbb{R}^{N}}f(u)u\,dx=\int_{\mathbb{R}^{N}}f(v)v\,dx,$
and hence
$0=M_{0}(u)=\Psi(u)-\int_{\mathbb{R}^{N}}H(u)\,dx\geq\Psi(v)-\int_{\mathbb{R}^{N}}H(v)\,dx=M_{0}(v).$
From Subsection˜4.1 and (4.2), there exists $\theta\in(0,1]$ such that $v_{\theta}\in{\mathcal{M}}_{0}$ and (F5) yields
$c_{0}\leq I(v_{\theta})=\frac{\theta^{N}}{N}\int_{\mathbb{R}^{N}}f(\theta v)\theta v\,dx\leq\frac{1}{N}\int_{\mathbb{R}^{N}}f(v)v\,dx=\frac{1}{N}\int_{\mathbb{R}^{N}}f(u)u\,dx=I(u)=c_{0}.$
Hence $\theta=1$ and $v\in{\mathcal{M}}_{0}$ is also a minimizer and this completes the proof. ∎
4.5. Sobolev-type inequality
Proof of Theorem˜1.4.
Consider the case $f(u)=|u|^{p-2}u$ with $p>2^{*}$. Fix any $u\in D(\Psi)\setminus\{0\}$. Then, from above considerations, it is clear that $I(m(u))\geq c_{0}$. Recalling that $m(u)=u_{\theta(u)}\in{\mathcal{M}}_{0}$, where $\theta=\theta(u)=\Theta(\Psi(u),u)$, we get
(4.15) $c_{0}\leq I(u_{\theta})=\frac{\theta^{N}}{N}\int_{\mathbb{R}^{N}}f\left(\theta u\right)\theta u\,dx=\frac{\theta^{N+p}}{N}\|u\|_{p}^{p}.$
Notice that the equality happens if and only if $m(u)=u_{\theta}$ is the minimizer on ${\mathcal{M}}_{0}$. Since $F(u)=\frac{1}{p}|u|^{p}$, $H(u)=\left(\frac{1}{p}+\frac{1}{N}\right)|u|^{p}=\frac{N+p}{Np}|u|^{p}$ and $M_{0}(m(u))=0$, it follows that
$\Psi(u)=\int_{\mathbb{R}^{N}}\frac{N+p}{Np}|\theta u|^{p}\,dx=\frac{N+p}{Np}\theta^{p}\int_{\mathbb{R}^{N}}|u|^{p}\,dx$
and that
$\theta=\left(\frac{Np}{N+p}\Psi(u)\right)^{1/p}\frac{1}{\|u\|_{p}}.$
Thus, from (4.15) we obtain
$c_{0}\leq\frac{\theta^{N+p}}{N}\|u\|_{p}^{p}=\frac{1}{N}\left(\frac{Np}{N+p}\Psi(u)\right)^{(N+p)/p}\frac{1}{\|u\|^{N}_{p}}$
and
$\left(\frac{N+p}{Np}\right)^{N+p}N^{p}c_{0}^{p}\|u\|_{p}^{Np}\leq\Psi(u)^{N+p}.$
Furthermore, the equality holds if and only if $m(u)=u_{\theta}$ is the minimizer of $\inf_{{\mathcal{M}}_{0}}I$, and the proof is completed. ∎
5. Positive mass case
In this section, the positive mass case is considered. Hereafter, $N\geq 1$ and (G1)–(G3) are assumed. For arguments which are similar to the zero mass case, we outline points where we need modifications from the zero mass case.
5.1. The projection into the Pohožaev set
We will show that there exists a continuous projection from $D(\Psi)\setminus\{0\}$ onto ${\mathcal{M}}_{1}$. To this end, we first prove
Lemma 5.1.
For every $\alpha>0$ and $u\in X_{1}\setminus\{0\}$ there is a unique $\Theta=\Theta(\alpha,u)>0$ such that
$\alpha=\int_{\mathbb{R}^{N}}-\left(\frac{1}{2}+\frac{1}{N}\right)\Theta^{2}u^{2}+G(\Theta u)+\frac{1}{N}g(\Theta u)\Theta u\,dx.$
Furthermore, the map $(0,\infty)\times X_{1}\setminus\{0\}\ni(\alpha,u)\mapsto\Theta(\alpha,u)\in(0,\infty)$ is of class ${\mathcal{C}}^{1}$.
Proof.
Set
${\mathcal{H}}(u,\theta)$ $=\int_{\mathbb{R}^{N}}-\left(\frac{1}{2}+\frac{1}{N}\right)\theta^{2}u^{2}+G(\theta u)+\frac{1}{N}g(\theta u)\theta u\,dx,$
$J(\alpha,u,\theta)$ $=\alpha-{\mathcal{H}}(u,\theta)\in{\mathcal{C}}^{1}\left((0,\infty)\times X_{1}\setminus\{0\}\times(0,\infty);\mathbb{R}\right).$
From $X_{1}\subset L^{\infty}(\mathbb{R}^{N})$, (G1) and (G3), it follows that for each fixed $u\in X_{1}\setminus\{0\}$,
(5.1) $\lim_{\theta\to 0}\left\|\frac{G(\theta u)}{\theta^{2}u^{2}}+\frac{1}{N}\frac{g(\theta u)}{\theta u}\right\|_{\infty}=0,\quad\lim_{\theta\to\infty}{\mathcal{H}}(u,\theta)=\infty.$
Furthermore, by (G2), for each $s\neq 0$,
(5.2) $G(s)=\int_{0}^{s}g(\tau)\,d\tau=\int_{0}^{s}\frac{g(\tau)}{\tau}\cdot\tau\,d\tau<\frac{g(s)}{s}\int_{0}^{s}\tau\,d\tau=\frac{1}{2}g(s)s.$
For any $u\in X_{1}\setminus\{0\}$ and $\theta\in(0,\infty)$ with ${\mathcal{H}}(u,\theta)\geq 0$, notice that (5.2) and (G2) yield
(5.3) $\frac{\partial}{\partial\theta}{\mathcal{H}}(u,\theta)$ $=\int_{\mathbb{R}^{N}}-2\left(\frac{1}{2}+\frac{1}{N}\right)\theta u^{2}+g(\theta u)u+\frac{1}{N}\left(g^{\prime}(\theta u)\theta u^{2}+g(\theta u)u\right)\,dx$
$=\theta^{-1}\left(2{\mathcal{H}}(u,\theta)+\int_{\mathbb{R}^{N}}-2G(\theta u)+\left(1-\frac{1}{N}\right)g(\theta u)\theta u+\frac{1}{N}g^{\prime}(\theta u)\theta^{2}u^{2}\,dx\right)$
$>\frac{\theta^{-1}}{N}\int_{\mathbb{R}^{N}}g^{\prime}(\theta u)\theta^{2}u^{2}-g(\theta u)\theta u\,dx\geq 0.$
From (5.1) and this fact, there exists a unique $\Theta_{1}=\Theta_{1}(u)>0$ such that ${\mathcal{H}}(u,\Theta_{1})=0$ and $(\Theta_{1},\infty)\ni\theta\mapsto{\mathcal{H}}(u,\theta)$ is strictly increasing. Therefore, for any given $\alpha>0$ and $u\in X_{1}\setminus\{0\}$, we may find a unique $\Theta=\Theta(\alpha,u)\in(\Theta_{1},\infty)$ such that $J(\alpha,u,\Theta)=\alpha-{\mathcal{H}}(u,\Theta)=0$. Moreover, by (5.3), we have
$\frac{\partial J}{\partial\theta}(\alpha,u,\Theta)<0.$
Hence, the regularity of $\Theta$ follows from the implicit function theorem. ∎
As in the zero mass case, consider $u_{\theta}:=\theta u(\cdot/\theta)$. Observe that
$\frac{d}{d\theta}I(u_{\theta})=N\theta^{N-1}\left(\Psi(u)-\int_{\mathbb{R}^{N}}-\left(\frac{1}{2}+\frac{1}{N}\right)\theta^{2}u^{2}+G(\theta u)+\frac{1}{N}g(\theta u)\theta u\,dx\right).$
By putting $\theta(u):=\Theta(\Psi(u),u)$, the following hold from Sections˜2 and 5.1:
Lemma 5.2.
For every $u\in D(\Psi)\setminus\{0\}$ there is a unique $\theta=\theta(u)>0$ such that $u_{\theta}\in{\mathcal{M}}_{1}$. Moreover the map
$m:D(\Psi)\setminus\{0\}\rightarrow{\mathcal{M}}_{1},\quad m(u):=u_{\theta(u)}$
is continuous where the topology on $D(\Psi)\setminus\{0\}$ is induced from $X_{1}$.
5.2. Minimizers on the Pohožaev set are critical points
For $u\in{\mathcal{M}}_{1}$, from $M_{1}(u)=0$ it follows that
(5.4) $I(u)=I(u)-\frac{N}{N+2}M_{1}(u)=\frac{2}{N+2}\left\{\Psi(u)+\int_{\mathbb{R}^{N}}\frac{1}{2}g(u)u-G(u)\,dx\right\}.$
From (5.2), the integrand of the right-hand side is nonnegative. Moreover, for each $s\in\mathbb{R}$ and $\theta>0$, (G2) implies
$\theta\frac{d}{d\theta}\left(\frac{1}{2}g(\theta u)\theta u-G(\theta u)\right)=\frac{1}{2}g^{\prime}(\theta u)(\theta u)^{2}-\frac{1}{2}g(\theta u)\theta u\geq 0.$
In particular, if $u\in D(\Psi)\setminus\{0\}$ and $\theta>0$, then
(5.5) $\frac{d}{d\theta}\int_{\mathbb{R}^{N}}\frac{1}{2}g(\theta u)\theta u-G(\theta u)\,dx\geq 0.$
To prove that minimizers on ${\mathcal{M}}_{1}$ are critical points of $I$, we need the following:
Lemma 5.3.
Let f(s):=s+g(s)f(s):=-s+g(s) and $u\in X_{1}\setminus\{0\}$ satisfy $\int_{\mathbb{R}^{N}}f(u)u\,dx>0$. Then for each $t>1$,
$\int_{\mathbb{R}^{N}}f(tu)tu\,dx>\int_{\mathbb{R}^{N}}f(u)u\,dx.$
Proof.
Denote $\beta:=\int_{\mathbb{R}^{N}}f(u)u\,dx>0$. For $t>1$ we observe that
$\frac{d}{dt}\int_{\mathbb{R}^{N}}f(tu)tu\,dx$ $=\int_{\mathbb{R}^{N}}-2tu^{2}+g^{\prime}(tu)tu^{2}+g(tu)u\,dx$
$=t\left(2\beta-2\int_{\mathbb{R}^{N}}g(u)u\,dx+\int_{\mathbb{R}^{N}}g^{\prime}(tu)u^{2}+\frac{g(tu)}{tu}u^{2}\,dx\right)$
$=t\left\{2\beta+\int_{\mathbb{R}^{N}}u^{2}\left(-2\frac{g(u)}{u}+g^{\prime}(tu)+\frac{g(tu)}{tu}\right)\,dx\right\}.$
Here (G2) and $g^{\prime}(s)s^{2}-g(s)s\geq 0$ imply
$\frac{d}{dt}\int_{\mathbb{R}^{N}}f(tu)tu\,dx\geq t\left\{2\beta+2\int_{\mathbb{R}^{N}}u^{2}\left(\frac{g(tu)}{tu}-\frac{g(u)}{u^{2}}\right)dx\right\}\geq 2\beta t>0$
and the proof is completed. ∎
We next define the mountain pass $c_{1,MP}$ by
$c_{1,MP}:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}I(\gamma(t)),\quad\Gamma:=\left\{\gamma\in{\mathcal{C}}([0,1];X_{1})\ :\ \gamma([0,1])\subset D(\Psi),\ \gamma(0)=0,\ I(\gamma(1))<0\right\}.$
Lemma 5.4.
For each $u\in{\mathcal{M}}_{1}$ there exists a path $\gamma_{u}\in{\mathcal{C}}([0,1];D(\Psi))$ such that $\max_{0\leq t\leq 1}I(\gamma_{u}(t))=I(u)$. Moreover, $c_{1}=c_{1,MP}>0$ holds.
Proof.
Let $u\in{\mathcal{M}}_{1}$ and consider $\gamma_{u}(t):=(u)_{Tt}=(Tt)u(\cdot/(Tt))$. It is immediate to see $\gamma\in{\mathcal{C}}([0,1];D(\Psi))$. Recalling the notation $\mathcal{H}$} in the proof of Subsection˜5.1, we see that
$\frac{d}{dt}I(\gamma_{u}(t))=NT(Tt)^{N-1}\left(\Psi(u)-{\mathcal{H}}(u,tT)\right).$
According to Subsection˜5.1 and $u\in{\mathcal{M}}_{1}$, for sufficiently large $T\gg 1$,
$\max_{0\leq t\leq 1}I(\gamma_{u}(t))=I(u),\quad I(\gamma_{u}(1))<0.$
Hence, the first assertion holds and this yields $c_{1,MP}\leq c_{1}$.
Next, we prove the inequality $c_{1,MP}>0$. If $|s|\ll 1$, then (G1) yields
$F(s)=-\frac{s^{2}}{2}+G(s)\leq-\frac{s^{2}}{2}+\frac{s^{2}}{4}=-\frac{1}{4}s^{2}.$
Thus, when $\|u\|_{X_{1}}\ll 1$, by $X_{1}\subset L^{\infty}(\mathbb{R}^{N})$, we have
$\Phi(u)\leq-\frac{1}{4}\|u\|_{2}^{2},\quad I(u)\geq b_{1}\|\nabla u\|_{2}^{2}+b_{2j_{0}}\|\nabla u\|_{2j_{0}}^{2j_{0}}+\frac{1}{4}\|u\|_{2}^{2}.$
Hence, $c_{1,MP}>0$ holds.
Finally, the inequality $c_{1}\leq c_{1,MP}$ will follow from $\gamma([0,1])\cap{\mathcal{M}}_{1}\neq\emptyset$ for any $\gamma\in\Gamma$. This assertion can be proved by the following argument. A similar argument to the above yields $M_{1}(u)>0$ for each $0<\|u\|_{X_{1}}\ll 1$. On the other hand, by
$M_{1}(\gamma(1))=I(\gamma(1))-\frac{1}{N}\Phi^{\prime}(\gamma(1))\gamma(1)$
and if $u\in D(\Psi)\setminus\{0\}$ satisfies $I(u)<0$, then (5.2) gives
$\Phi^{\prime}(u)u=\int_{\mathbb{R}^{n}}f(u)u\,dx=\int_{\mathbb{R}^{N}}-u^{2}+g(u)u\,dx\geq 2\int_{\mathbb{R}^{N}}-\frac{u^{2}}{2}+G(u)\,dx=2\left(-I(u)+\Psi(u)\right)>0.$
Hence, $M_{1}(\gamma(1))<0$ and there exists $\tau\in(0,1)$ such that $M_{1}(\gamma(\tau_{1}))=0$. This completes the proof. ∎
Lemma 5.5.
Let $u_{0}\in{\mathcal{M}}_{1}$ satisfy $I(u_{0})=c_{1}$. Then $u_{0}$ is a critical point of $I$.
Proof.
Set
$H(s):=F(s)+\frac{1}{N}f(s)s=-\left(\frac{1}{2}+\frac{1}{N}\right)s^{2}+G(s)+\frac{1}{N}g(s)s.$
Then the proof of Subsection˜4.3 works until (4.9). After (4.9), we need to prove $\theta_{n}<1$ under $\Psi(v_{n})<\Psi(u_{0})$. Recall the notation $\mathcal{H}$} in the proof of Subsection˜5.1. Since $u_{0}\in{\mathcal{M}}_{1}$ gives
${\mathcal{H}}(u_{0},1)=\int_{\mathbb{R}^{N}}H(u_{0})\,dx=\Psi(u_{0})>0,$
(5.3) and (G2) lead to
$\int_{\mathbb{R}^{N}}H^{\prime}(u_{0})u_{0}\,dx=\frac{\partial{\mathcal{H}}}{\partial\theta}(u_{0},1)>\frac{1}{N}\int_{\mathbb{R}^{N}}g^{\prime}(u_{0})u_{0}^{2}-g(u_{0})u_{0}\,dx\geq 0.$
Hence, the argument for proving $\theta_{n}<1$ also works in the positive mass case.
Finally, if $\theta_{n}<1$ and $w_{n}=\theta_{n}v_{t_{n}}(x/\theta_{n})$, thanks to $0<c_{1}=I(u_{0})$ and Subsection˜5.2,
$I(u_{0})\leq I(w_{n})=\frac{\theta_{n}^{N}}{N}\int_{\mathbb{R}^{N}}f(\theta_{n}v_{t_{n}})\theta_{n}v_{t_{n}}\,dx$ $\leq\frac{\theta_{n}^{N}}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx$
$=\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx-\frac{(1-\theta_{n}^{N})}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx.$
By $0<c_{1}=I(u_{0})=\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx/N$, we can repeat the same reasoning as for the zero mass case in Subsection˜4.3 to show that every minimizer of $I$ on ${\mathcal{M}}_{1}$ is a critical point. ∎
5.3. Proof of Theorem˜1.6
Proof of Theorem˜1.6.
Remark that as in the proof of Subsection˜5.2, for a minimzing sequence $(u_{n})_{n\in\mathbb{N}}\subset{\mathcal{M}}_{1}$, by $M_{1}(u_{n})=0$, we obtain
$NI(u_{n})=\Phi^{\prime}(u_{n})u_{n}=\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx\geq 2\int_{\mathbb{R}^{N}}F(u_{n})\,dx=2\left(-I(u_{n})+\Psi(u_{n})\right).$
Hence, $(\Psi(u_{n}))_{n\in\mathbb{N}}$, $(\Phi^{\prime}(u_{n})u_{n})_{n\in\mathbb{N}}$ and $(\Phi(u_{n}))_{n\in\mathbb{N}}$ are bounded according to $I(u_{n})=c_{1}+o(1)$. When $N\geq 3$, by $X_{1}\subset X_{0}\subset L^{2^{*}}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$, $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{2^{*}}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$. From (G1), we may find $C>0$, which depends on $\sup_{n}\|u_{n}\|_{\infty}$ such that
$G(u_{n}(x))\leq\frac{1}{4}u_{n}^{2}(x)+C|u_{n}(x)|^{2^{*}}\quad\text{for all $n$ and $x\in\mathbb{R}^{N}$}.$
Hence,
$c_{1}+o(1)=I(u_{n})\geq\frac{1}{2}\|u_{n}\|_{2}^{2}-\int_{\mathbb{R}^{N}}G(u_{n})\,dx\geq\frac{1}{4}\|u_{n}\|_{2}^{2}-C\|u_{n}\|_{2^{*}}^{2^{*}}\geq\frac{1}{4}\|u_{n}\|_{2}^{2}-C\|\nabla u_{n}\|_{2}^{2^{*}}.$
By $\|\nabla u_{n}\|_{2}^{2}/2+b_{2j_{0}}\|\nabla u_{n}\|_{2j_{0}}^{2j_{0}}\leq\Psi(u_{n})$, $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{2}(\mathbb{R}^{N})$, hence in $X_{1}$.
On the other hand, when $N=1,2$, we claim that (G2) and (G3) imply that
(5.6) $\lim_{|s|\to\infty}\left(g(s)s-2G(s)\right)=\infty.$
In fact, by (G3), we may choose $(s_{n})_{n\in\mathbb{N}}$ so that
$s_{n}<s_{n+1},\quad s_{n}\to\infty,\quad\frac{g(s_{n})}{s_{n}}+1\leq\frac{g(s)}{s}\quad\text{for $s\geq s_{n+1}$}.$
Hence, if $s\geq s_{n+1}$, then
$2G(s)$ $=2\int_{0}^{s_{n}}\frac{g(\tau)}{\tau}\tau\,d\tau+2\int_{s_{n}}^{s}\frac{g(\tau)}{\tau}\tau\,d\tau$
$\leq 2\frac{g(s_{n})}{s_{n}}\int_{0}^{s_{n}}\tau\,d\tau+2\frac{g(s)}{s}\int_{s_{n}}^{s}\tau\,d\tau$
$=\left(\frac{g(s_{n})}{s_{n}}-\frac{g(s)}{s}\right)s_{n}^{2}+g(s)s.$
The choice of $(s_{n})_{n\in\mathbb{N}}$ and $s\geq s_{n+1}$ lead to
$g(s)s-2G(s)\geq\left(\frac{g(s)}{s}-\frac{g(s_{n})}{s_{n}}\right)s_{n}^{2}\geq s_{n}^{2},$
which implies $g(s)s-2G(s)\to\infty$ as $s\to\infty$. Similarly, we may verify $g(s)s-2G(s)\to\infty$ as $s\to$-$\infty$ and (5.6) holds.
Recalling that
$\int_{\mathbb{R}^{N}}g(u_{n})u_{n}-2G(u_{n})\,dx=\Phi^{\prime}(u_{n})u_{n}-2\Phi(u_{n})$
is bounded and that each $u_{n}$ is $1$-Lipschitz in view of $u_{n}\in D(\Psi)$, we observe from (5.2) and (5.6) that $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{\infty}(\mathbb{R}^{N})$. Furthermore, by (5.2) and (5.6), for each $t>0$ there exists $\delta_{t}>0$ such that $g(s)s-2G(s)\geq\delta_{t}$ if $|s|\geq t$. Therefore, $(|\{|u_{n}|\geq t\}|)_{n\in\mathbb{N}}$ is also bounded where $|A|$ denotes the $N$-dimensional Lebesgue measure of $A\subset\mathbb{R}^{N}$. By (G1), choose $t_{0}>0$ so that $|G(t)|\leq t^{2}/4$ for each $|t|\leq t_{0}$. Then
$c_{1}+o(1)=I(u_{n})$ $\geq\frac{1}{2}\|u_{n}\|_{2}^{2}-\int_{\mathbb{R}^{N}}G(u_{n})\,dx$
$\geq\frac{1}{4}\|u_{n}\|_{2}^{2}-\int_{|u_{n}|\geq t_{0}}G(u_{n})\,dx$
$\geq\frac{1}{4}\|u_{n}\|_{2}^{2}-\left(\max_{|t|\leq\sup_{n}\|u_{n}\|_{\infty}}G(t)\right)\left|\left\{|u_{n}|\geq t_{0}\right\}\right|$
and $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{2}(\mathbb{R}^{N})$. Thus, when $N=1,2$, $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X_{1}$.
Now we use Section˜3 with $a=(N+2)/(2N)$. Then we obtain $K\in\{0,1,2,\dots\}\cup\{\infty\}$, $(y_{n}^{j})_{n}$, $\widetilde{u}_{j}$ for $0\leq j<K+1$ such that
(5.7) $u_{n}(\cdot+y_{n}^{j})\rightharpoonup\widetilde{u}_{j},\quad u_{n}(x+y_{n}^{j})\rightharpoonup\widetilde{u}_{j}(x)\quad\text{a.e. $\mathbb{R}^{N}$},$
$\lim_{n\to\infty}\left(\Psi(u_{n})+a\|u_{n}\|_{2}^{2}\right)\geq\sum_{j=0}^{K}\left(\Psi(\widetilde{u}_{j})+a\|\widetilde{u}_{j}\|_{2}^{2}\right),$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}G(u_{n})\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}G(\widetilde{u}_{j})\,dx,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}g(u_{n})u_{n}\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}g(\widetilde{u}_{j})\widetilde{u}_{j}\,dx.$
Since $M_{1}(u_{n})=0$, (5.7) with $a=(N+2)/(2N)$ yields
(5.8) $0=\lim_{n\to\infty}M_{1}(u_{n})\geq\sum_{j=0}^{K}M_{1}(\widetilde{u}_{j}).$
When $K=0$ and $\widetilde{u}_{0}=0$, from $\int_{\mathbb{R}^{N}}G(u_{n})\,dx=o(1)=\int_{\mathbb{R}^{N}}g(u_{n})u_{n}\,dx$ and (5.8), it follows that $\Psi(u_{n})+(N+2)\|u_{n}\|_{2}^{2}/(2N)\to 0$, however, this gives the following contradiction $0<c_{1}=\lim_{n\to\infty}I(u_{n})=0$. Hence, either $K\geq 1$ or $K=0$ with $\widetilde{u}_{0}\neq 0$.
As in the proof of Theorem˜1.2, set $J:=\left\{j\ :\ \widetilde{u}_{j}\neq 0\right\}$. Then $J\neq\emptyset$ and (5.8) becomes
(5.9) $0\geq\sum_{j\in J}M_{1}(\widetilde{u}_{j}).$
Therefore, there exists $j_{0}\in J$ such that $M_{1}(\widetilde{u}_{j_{0}})\leq 0$. By $\widetilde{u}_{j_{0}}\neq 0$, we claim $\theta_{j_{0}}=\Theta(\Psi(\widetilde{u}_{j_{0}}),\widetilde{u}_{j_{0}})\leq 1$. Indeed, by recalling $H$ and $\mathcal{H}$} in the proofs of Subsections˜5.1 and 5.2, $\widetilde{u}_{j_{0}}\neq 0$ and $M_{1}(\widetilde{u}_{j_{0}})\leq 0$, we see that
$0<\Psi(\widetilde{u}_{j_{0}})\leq\int_{\mathbb{R}^{N}}H(\widetilde{u}_{j_{0}})\,dx={\mathcal{H}}(\widetilde{u}_{j_{0}},1).$
Therefore, (5.3) leads to
$\frac{d}{d\theta}\int_{\mathbb{R}^{N}}H(\theta\widetilde{u}_{j_{0}})\,dx\big|_{\theta=1}=\frac{\partial}{\partial\theta}{\mathcal{H}}(\widetilde{u}_{j_{0}},\theta)\big|_{\theta=1}>\frac{1}{N}\int_{\mathbb{R}^{N}}g^{\prime}(u)u^{2}-g(u)u\,dx\geq 0.$
Using (5.3) again, we observe that the map $1,\infty)\ni\theta\mapsto\int_{\mathbb{R}^{N}}H(\theta\widetilde{u}_{j_{0}})\,dx$ is strictly increasing. Since $\Theta(\Psi(\widetilde{u}_{j_{0}}),\widetilde{u}_{j_{0}})>0$ is the unique solution to
$\Psi(\widetilde{u}_{j_{0}})=\int_{\mathbb{R}^{N}}H(\theta\widetilde{u}_{j_{0}})\,dx,$
we get $\theta_{j_{0}}\leq 1$. Thus by writing $h(s):=g(s)s/2-G(s)\geq 0$, from $(\widetilde{u}_{j_{0}})_{\theta_{j_{0}}}\in{\mathcal{M}}_{1}$, (5.4), (5.5), the weak lower semicontinuity of $\Psi$ and Fatou’s lemma, it follows that
$c_{1}\leq I((\widetilde{u}_{j_{0}})_{\theta_{j_{0}}})$ $=\frac{2\theta_{j_{0}}^{N}}{N+2}\left\{\Psi(\widetilde{u}_{j_{0}})+\int_{\mathbb{R}^{N}}h(\theta_{j_{0}}\widetilde{u}_{j_{0}})\,dx\right\}$
$\leq\frac{2}{N+2}\left\{\Psi(\widetilde{u}_{j_{0}})+\int_{\mathbb{R}^{N}}h(\widetilde{u}_{j_{0}})\,dx\right\}$
$\leq\frac{2}{N+2}\liminf_{n\to\infty}\left\{\Psi(u_{n}(\cdot+y^{j_{0}}_{n}))+\int_{\mathbb{R}^{N}}h(u_{n}(\cdot+y^{j_{0}}_{n}))\,dx\right\}$
$=\liminf_{n\to\infty}I(u_{n})=c_{1}.$
Therefore, $\theta_{j_{0}}=1$ and $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{1}$ is a minimizer. Furthermore, by Subsection˜5.2, $\widetilde{u}_{j_{0}}$ is a critical point of $I$.
We next prove $\|u_{n}(\cdot+y^{j_{0}}_{n})-\widetilde{u}_{j_{0}}\|_{X_{1}}\to 0$. To this end, we first show $\sharp J=1$. If $\sharp J\geq 2$, then since $M_{1}(\widetilde{u}_{j_{0}})=0$, (5.9) becomes
$0\geq\sum_{j\in J\setminus\{j_{0}\}}M_{1}(\widetilde{u}_{j}).$
Thus, there exists $j_{1}\in J\setminus\{j_{0}\}$ such that $M_{1}(\widetilde{u}_{j_{1}})\leq 0$ and we can repeat the above argument to obtain
$\widetilde{u}_{j_{1}}\in{\mathcal{M}}_{1},\quad c_{1}=I(\widetilde{u}_{j_{1}})=\frac{2}{N+2}\left(\Psi(\widetilde{u}_{j_{1}})+\int_{\mathbb{R}^{N}}h(\widetilde{u}_{j_{1}})\,dx\right).$
Choose $R_{0}\gg 1$ so large that
$\frac{c_{1}}{2}<\frac{2}{N+2}\int_{B(0,R_{0})}1-\sqrt{1-|\nabla\widetilde{u}_{j_{k}}|^{2}}+h(\widetilde{u}_{j_{k}})\,dx=:K_{B(0,R_{0})}(\widetilde{u}_{j_{k}})\quad(k=0,1).$
By recalling $|y^{j_{0}}_{n}-y^{j_{1}}_{n}|\to\infty$ and $h(s)\geq 0$ for each $s\in\mathbb{R}$, the weak lower semicontinuity of $\Psi_{B(y,r)}$ and Fatou’s lemma give
$c_{1}$ $<\sum_{k=0}^{1}K_{B(0,R_{0})}\left(\widetilde{u}_{j_{k}}\right)$
$\leq\liminf_{n\to\infty}\sum_{k=0}^{1}K_{B(y^{j_{k}}_{n},R_{0})}\left(u_{n}\right)$
$\leq\liminf_{n\to\infty}\frac{2}{N+2}\int_{\mathbb{R}^{N}}1-\sqrt{1-|\nabla u_{n}|^{2}}+h(u_{n})\,dx=\liminf_{n\to\infty}I(u_{n})=c_{1},$
which is a contradiction. Hence, $\sharp J=1$. From (5.7) with $a=(N+2)/(2N)$, (5.8), $\sharp J=1$ and $M_{1}(\widetilde{u}_{j_{0}})=0$, we infer that
$\Psi(u_{n}(\cdot+y^{j_{0}}_{n}))+\frac{N+2}{2N}\|u_{n}(\cdot+y^{j_{0}}_{n})\|_{2}^{2}\to\Psi(\widetilde{u}_{j_{0}})+\frac{N+2}{2N}\|\widetilde{u}_{j_{0}}\|_{2}^{2}.$
Combining this with Section˜2, we have $\|u_{n}(\cdot+y^{j_{0}}_{n})-\widetilde{u}_{j_{0}}\|_{X_{1}}\to 0$.
We move to the proof that each minimizer has a constant sign and an argument is similar to the above. Let $u\in{\mathcal{M}}_{1}$ be any minimizer of $c_{1}$ and suppose $u_{+},u_{-}\not\equiv 0$. Since $0=M_{1}(u)=M_{1}(u^{+})+M_{1}(u^{-})$, either $M_{1}(u_{+})\leq 0$ or $M_{1}(u_{-})\leq 0$. Suppose $M_{1}(u_{+})\leq 0$. By $u_{+}\not\equiv 0$, there exists $\theta_{+}\leq 1$ such that $(u_{+})_{\theta_{+}}\in{\mathcal{M}}_{1}$. From (5.5), $h(s)=g(s)s/2-G(s)\geq 0$, $\Psi(u^{-})>0$ and
$\Psi(u)=\Psi(u_{+})+\Psi(u_{-}),\quad\int_{\mathbb{R}^{N}}h(u)u\,dx=\int_{\mathbb{R}^{N}}h(u_{+})u_{+}\,dx+\int_{\mathbb{R}^{N}}h(u_{-})u_{-}\,dx,$
it follows that
$c_{1}\leq I((u_{+})_{\theta_{+}})$ $=\frac{2\theta_{+}^{N}}{N+2}\left\{\Psi(u_{+})+\int_{\mathbb{R}^{N}}h(\theta_{+}u_{+})\,dx\right\}$
$\leq\frac{2}{N+2}\left\{\Psi(u_{+})+\int_{\mathbb{R}^{N}}h(u_{+})\,dx\right\}$
$<\frac{2}{N+2}\left\{\Psi(u)+\int_{\mathbb{R}^{N}}h(u)\,dx\right\}=I(u)=c_{1},$
which is a contradiction. A similar contradiction is obtained if $M_{1}(u^{-})\leq 0$ holds. Therefore, either $u_{+}\equiv 0$ or $u_{-}\equiv 0$ holds, and the strict sign property follows from the same reasoning as in the zero mass case.
The existence of a radial minimizer follows as in the proof of Theorem˜1.2. Indeed, let $u\in{\mathcal{M}}_{1}$ be any minimizer and write $v:=u^{*}$ when $u$ is positive, and $v:=-u^{*}$ when $u$ is negative. Then we may verify $M_{1}(v)\leq 0$, by Subsection˜5.1, there exists $\theta\in(0,1]$ such that $v_{\theta}\in{\mathcal{M}}_{1}$. Then we may infer from Subsection˜5.2 and $I(w)=\int_{\mathbb{R}^{N}}f(w)w\,dx/N$ for any $w\in{\mathcal{M}}_{1}$ that
$0<c_{1}\leq I(v_{\theta})=\frac{\theta^{N}}{N}\int_{\mathbb{R}^{N}}f(\theta v)\theta v\,dx\leq\frac{1}{N}\int_{\mathbb{R}^{N}}f(v)v\,dx=\frac{1}{N}\int_{\mathbb{R}^{N}}f(u)u\,dx=I(u)=c_{1}.$
Thus, $\theta=1$ and $v\in{\mathcal{M}}_{1}$ is a radial minimizer. ∎
6. Nonradial solutions
In this section, Theorem˜1.7 is proved. Throughout this section, we suppose $N\geq 4$ and fix any $k_{1},k_{2}\in\mathbb{N}$ so that $k_{1},k_{2}\geq 2$ and $N-k_{1}-k_{2}\geq 0$ and consider
${\mathcal{O}}={\mathcal{O}}(k_{1})\times{\mathcal{O}}(k_{2})\times\{\mathrm{id}\}\subset{\mathcal{O}}(N).$
By writing $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\mathbb{R}^{N-k_{1}-k_{2}}$ (when $N=k_{1}+k_{2}$, we write $x=(x_{1},x_{2})$), recall $X_{\mathcal{O}}$, $\tau$ and $X_{\tau}$:
$X_{\mathcal{O}}:=\left\{u\in X\ :\ u(\rho x)=u(x)\quad\text{for all $x\in\mathbb{R}^{N}$ and $\rho\in{\mathcal{O}}$}\right\}$
and
$(\tau u)(|x_{1}|,|x_{2}|,x_{3}):=-u(|x_{2}|,|x_{1}|,x_{3}),\quad X_{\tau}:=\left\{u\in X_{\mathcal{O}}\ :\ \tau u=u\right\},$
where $X=X_{0}$ in the zero mass case and $X=X_{1}$ in the positive mass case. Since $f$ (resp. $g$) is odd, $\Psi(-u)=\Psi(u)$ and $\Phi(-u)=\Phi(u)$. According to the principle of symmetric criticality (see [27, 18]), any critical point of $I$ in $X_{\tau}$ is also a critical point in $X_{\mathcal{O}}$, and by applying the principle again, it is a critical point in $X$. Thus, it suffices to find a critical point of $I|_{X_{\tau}}$.
We first prove that for the profile decomposition of $(u_{n})_{n\in\mathbb{N}}$ in $X_{\mathcal{O}}$, we may assume $y_{n}^{j}\in\{0\}\times\{0\}\times\mathbb{R}^{N-k_{1}-k_{2}}$:
Lemma 6.1.
Let $(u_{n})_{n\in\mathbb{N}}\subset X_{\mathcal{O}}$ be bounded and $K\in\{0\}\cup\mathbb{N}\cup\{\infty\}$, $(y_{n}^{j})_{n\in\mathbb{N}}\subset\mathbb{R}^{N}$, $\widetilde{u}_{j}\in X$ for $0\leq j<K+1$ be as in Section˜3. Then for each $j$, $\widetilde{u}_{j}\in X_{\mathcal{O}}$ and
$\limsup_{n\to\infty}\left\{|y_{n,1}^{j}|+|y_{n,2}^{j}|\right\}<\infty.$
In particular, we may suppose $y_{n,1}^{j}=0$ and $y_{n,2}^{j}=0$ for any $1\leq j<K+1$ and $n$.
Proof.
The assertion $\widetilde{u}_{j}\in X_{\mathcal{O}}$ follows from the closedness of $X_{{\mathcal{O}}}$ in $X$. The latter assertion may be proved as in [34]. Let $j\in\mathbb{N}$. By $u_{n}(\cdot+y_{n}^{j})\rightharpoonup\widetilde{u}_{j}\not\equiv 0$ weakly in $X$ and $u_{n}(\cdot+y_{n}^{j})\to\widetilde{u}_{j}$ in $L^{p_{0}}_{\rm loc}(\mathbb{R}^{N})$, there exist $r_{j}>0$ and $N_{j}\in\mathbb{N}$ such that for all $n\geq N_{j}$, $\|u_{n}\|_{L^{p_{0}}(B_{r_{j}}(y_{n}^{j}))}\geq\|\widetilde{u}_{j}\|_{L^{p_{0}}(B_{r_{j}}(0))}/2>0$. Since $k_{1},k_{2}\geq 2$, one may find $m_{1}=m_{1}(k_{1},|y_{n,1}^{j}|)$ (resp. $m_{2}=m_{2}(k_{2},|y_{n,2}^{j}|)$) disjoint balls $B_{1},\dots,B_{m_{1}}$ (resp. $B_{1},\dots,B_{m_{2}}$) with radius $r_{j}$ and centered on $\partial B_{|y_{n,1}^{j}|}(0)\cap\mathbb{R}^{k_{1}}\times\{0\}\times\{0\}$ (resp. $\partial B_{|y_{n,1}^{j}|}(0)\cap\{0\}\times\mathbb{R}^{k_{2}}\times\{0\}$), where $m(k_{i},r)\to\infty$ as $r\to\infty$ for $i=1,2$. Since $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X$ and $u_{n}\in X_{{\mathcal{O}}}$, for each $n\geq N_{j}$ and $i=1,2$,
$\infty>\sup_{k\in\mathbb{N}}\|u_{k}\|_{p_{0}}^{p_{0}}\geq\sum_{\ell=1}^{m_{i}}\|u_{n}\|_{L^{p_{0}}(B_{\ell})}^{p_{0}}=m_{i}\|u_{n}\|_{L^{p_{0}}(B_{r_{j}}(0))}\geq\frac{m_{i}}{2}\|\widetilde{u}_{j}\|_{L^{p_{0}}(B_{r_{j}}(0))}>0.$
Since $m_{i}=m_{i}(k_{i},r)\to\infty$ as $r\to\infty$, $(y^{j}_{n,i})_{n\in\mathbb{N}}$ $(i=1,2)$ are bounded (the bound depends on $\|\widetilde{u}_{j}\|_{L^{p_{0}}(B_{r_{j}}(0))}$ ) and we complete the proof. ∎
Proof of Theorem˜1.7.
We first prove the existence of minimizer corresponding to $\inf_{{\mathcal{M}}_{\tau}}I$, where ${\mathcal{M}}_{\tau}:={\mathcal{M}}\cap X_{\tau}$. Let $(u_{n})_{n\in\mathbb{N}}\subset{\mathcal{M}}_{\tau}$ satisfy $I(u_{n})\to\inf_{{\mathcal{M}}_{\tau}}I\in[\inf_{{\mathcal{M}}}I,\infty)$. As in the proofs of Theorems˜1.2 and 1.6, $(u_{n})_{n\in\mathbb{N}}$ is bounded in $X$. By Section˜3 and Section˜6, there exist $K\in\{0\}\cup\mathbb{N}\cup\{\infty\}$, $(y^{j}_{n})_{n\in\mathbb{N}}\subset\{0\}\times\{0\}\times\mathbb{R}^{N-k_{1}-k_{2}}$ and $\widetilde{u}_{j}$ such that
$u_{n}(\cdot+y_{n}^{j})\rightharpoonup\widetilde{u}_{j},\quad u_{n}(x+y_{n}^{j})\to\widetilde{u}_{j}(x)\quad\text{a.e. $\mathbb{R}^{N}$},$
$\lim_{n\to\infty}\Psi(u_{n})\geq\sum_{j=0}^{K}\Psi_{0}(\widetilde{u}_{j})\quad\left(\text{resp.}\ \lim_{n\to\infty}\left\{\Psi(u_{n})+a\left\|u_{n}\right\|^{2}_{2}\right\}\geq\sum_{j=0}^{K}\left\{\Psi(\widetilde{u}_{j})+a\|\widetilde{u}_{j}\|_{2}^{2}\right\}\right)$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}F(u_{n})\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}F(\widetilde{u}_{j})\,dx\quad\left(\text{resp.}\ \lim_{n\to\infty}\int_{\mathbb{R}^{N}}G(u_{n})\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}G(\widetilde{u}_{j})\,dx\right),$
$\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j})\widetilde{u}_{j}\,dx\quad\left(\text{resp.}\ \lim_{n\to\infty}\int_{\mathbb{R}^{N}}g(u_{n})u_{n}\,dx=\sum_{j=0}^{K}\int_{\mathbb{R}^{N}}g(\widetilde{u}_{j})\widetilde{u}_{j}\,dx\right).$
Then as in the proof of Theorems˜1.2 and 1.6, we may prove that $\inf_{{\mathcal{M}}_{\tau}}I$ is attained.
Let $u\in{\mathcal{M}}_{\tau}$ satisfy $I(u)=\inf_{{\mathcal{M}}_{\tau}}I$. Then we may use the argument in the proof of Subsections˜4.3 and 5.2 to verify that $u$ is a critical point of $I$ on $X_{\tau}$. As pointed in the beginning of this section, $u$ is a critical point of $I$ in $X$, and hence Section˜2 implies that $u$ is a classical solution of (1.1).
When $k_{1}=k_{2}$, we next prove $I(u)\geq 2\inf_{{\mathcal{M}}}I$ where $u\in{\mathcal{M}}_{\tau}$ is a minimizer corresponding to $\inf_{{\mathcal{M}}_{\tau}}I$. Set
$\Omega_{1}$ $=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{N}\ :\ |x_{1}|>|x_{2}|\},$
$\Omega_{2}$ $=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{N}\ :\ |x_{1}|<|x_{2}|\}.$
In this case, $\tau$ is identified with $\tau(x_{1},x_{2},x_{3})=(x_{2},x_{1},x_{3})$ and notice that $\tau\Omega_{1}=\Omega_{2}$ and $|\mathbb{R}^{N}\setminus(\Omega_{1}\cup\Omega_{2})|=0$. Since $u\in X_{\tau}\setminus\{0\}$ and $u=0$ on $\partial\Omega_{i}$ $(i=1,2)$, we get $\bm{1}_{\Omega_{i}}u\in X$ ($i=1,2$) and $0=M(u)=2M(\bm{1}_{\Omega_{i}}u)$. Thus, $\bm{1}_{\Omega_{i}}u\in{\mathcal{M}}$ and
$I(u)=I(\bm{1}_{\Omega_{1}}u)+I(\bm{1}_{\Omega_{2}}u)=2I(\bm{1}_{\Omega_{1}}u)\geq 2\inf_{{\mathcal{M}}}I.$
Finally, we show $I(u)>2\inf_{\mathcal{M}}I$. From $\bm{1}_{\Omega_{1}}u\in{\mathcal{M}}$, if $I(\bm{1}_{\Omega_{1}}u)=\inf_{{\mathcal{M}}}I$, then by Theorems˜1.2 and 1.6, either $\bm{1}_{\Omega_{1}}u>0$ or else $\bm{1}_{\Omega_{1}}u<0$. However, this is a contradiction. Thus, $I(\bm{1}_{\Omega_{1}}u)>\inf_{{\mathcal{M}}}I$ and $I(u)=2I(\bm{1}_{\Omega_{1}}u)>2\inf_{{\mathcal{M}}}I$ holds. ∎
Acknowledgements
B. Bieganowski was partly supported by the National Science Centre, Poland (grant no. 2023/51/B/ST1/00968). J. Mederski was partly supported by the National Science Centre, Poland (grant no. 2025/57/B/ST1/04804). This work was supported by JSPS KAKENHI Grant Number JP24K06802. This work was initiated and partially done when N. Ikoma visited Warsaw and he would like express his gratitude to University of Warsaw and Institute of Mathematics of the Polish Academy of Sciences.

### v2

@@
 $\lim_{h\to 0}\frac{I(u+t\varphi)-I(u)}{h}=\int_{\mathbb{R}^{N}}\frac{\nabla u\cdot\nabla\varphi}{\sqrt{1-|\nabla u|^{2}}}-f(u)\varphi\,dx$
-and $u$ is a weak solution of (1.1) provided $u$ is a critical point of $I$ with $\|\nabla u\|_{\infty}<1$.
+and $u$ is a weak solution of (1.1) provided $u$ is a critical point of II with $\|\nabla u\|_{\infty}<1$.
 However, it should be pointed out that for the well-definedness of $\Psi$, the condition $|\nabla u|\leq 1$ for a.e.
 $x\in\mathbb{R}^{N}$ is required and it is not possible to find a function space on which $\Psi$ becomes of class ${\mathcal{C}}^{1}$.
-Therefore, we need to go beyond the smooth critical point theory to treat $I$ directly on some function space, and the nonsmoothness of $\Psi$ makes the problem complicated.
+Therefore, we need to go beyond the smooth critical point theory to treat II directly on some function space, and the nonsmoothness of $\Psi$ makes the problem complicated.
 The problem
@@
 We start our considerations with the zero mass case in the spirit of Berestycki and Lions [5].
-1.1. The zero mass case
+$1.$ The zero mass case
 In this subsection, the results for the zero mass case are stated.
@@
     $|F(s)|\lesssim|f(s)s|$ for $s\in\mathbb{R}$, that is, there exists a $C\geq 0$ such that $|F(s)|\leq C|f(s)s|$ for every $s\in\mathbb{R}$.
-Remark 1.1.
+Remark $1.$
   1. (i)
@@
 Since $f(s)=2k_{0}s^{2k_{0}-1}/(1+s^{2k_{0}})$, it is immediate to verify (F1)–(F5), however, (F6) is not satisfied.
-By (1.4), (F1) and (F2), we may verify $\Phi\in{\mathcal{C}}^{1}(X_{0};\mathbb{R})$ and it allows us to define the notion of a critical point of $I$ as in [32].
-Namely, $u\in X_{0}$ is said to be a critical point of $I$ if
+By (1.4), (F1) and (F2), we may verify $\Phi\in{\mathcal{C}}^{1}(X_{0};\mathbb{R})$ and it allows us to define the notion of a critical point of II as in [32].
+Namely, $u\in X_{0}$ is said to be a critical point of II if
 (1.5) $\Psi(v)-\Psi(u)-\Phi^{\prime}(u)(v-u)\geq 0\quad\text{for all $v\in X_{0}$},$
@@
 $\|u\|_{X_{1}}:=\|u\|_{H^{1}}+\|\nabla u\|_{2j_{0}}.$
-Then, the variational functional $I$ is given by (1.2), $\Psi$ is defined as before with the natural domain
+Then, the variational functional II is given by (1.2), $\Psi$ is defined as before with the natural domain
 $D(\Psi):=\left\{u\in X_{1}\ :\ \|\nabla u\|_{\infty}\leq 1\right\}$
@@
 Remark that $\Psi$ is a convex functional on a convex domain $D(\Psi)$ and $\Phi\in{\mathcal{C}}^{1}(X_{1};\mathbb{R})$.
-Therefore, the notation of critical points of $I$ on $X_{1}$ is defined as in (1.5) by changing $X_{0}$ to $X_{1}$.
+Therefore, the notation of critical points of II on $X_{1}$ is defined as in (1.5) by changing $X_{0}$ to $X_{1}$.
 We also introduce the Pohožaev set as
@@
 This detailed analysis also allows us to prove that every minimizer has a constant sign, and this fact leads to the existence of radial minimizers.
-Even though we know the existence of minimizers, it is worth noting that whether these minimizers become critical points of $I$ is not trivial or immediate, and this point is more delicate than the semilinear case.
+Even though we know the existence of minimizers, it is worth noting that whether these minimizers become critical points of II is not trivial or immediate, and this point is more delicate than the semilinear case.
 For the semilinear case, a similar assertion is proved in [29, Lemma 3.3] where the differentiability of functional originated from the differential operator is utilized.
 In our case, the corresponding functional is $\Psi$, which is not smooth on $X_{0}$ or $X_{1}$.
-In addition, the nonsmoothness of $M_{0}$ and $M_{1}$ affect the properties of the projections to ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$, and the analysis of the projections onto ${\mathcal{M}}_{0}$, ${\mathcal{M}}_{1}$ and functional $I$ is more delicate and complicated.
+In addition, the nonsmoothness of $M_{0}$ and $M_{1}$ affect the properties of the projections to ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$, and the analysis of the projections onto ${\mathcal{M}}_{0}$, ${\mathcal{M}}_{1}$ and functional II is more delicate and complicated.
 These issues will be solved in Subsections˜4.3 and 5.2 and it is one of novelties in this paper.
@@
 $M_{1}(u)=0$) is natural in the sense that a critical point of $I|_{{\mathcal{M}}_{0}}$ (resp.
-$I|_{{\mathcal{M}}_{1}}$) provides a critical point of $I$.
+$I|_{{\mathcal{M}}_{1}}$) provides a critical point of II.
 Therefore, it is interesting to see whether the same is true for higher energies or other nonsmooth constraints.
 The structure of the paper is as follows.
-In Section˜2, the functional setting is recalled and some preliminary properties related to the functionals $I$ and $\Psi$ are proved.
+In Section˜2, the functional setting is recalled and some preliminary properties related to the functionals II and $\Psi$ are proved.
 In Section˜3 we develop profile decompositions for bounded sequences in $X_{0}$ and $X_{1}$.
@@
 $M(u):=\begin{dcases}M_{0}(u)&\text{(the zero mass case)},\\ M_{1}(u)&\text{(the positive mass case)}.\end{dcases}$
-Finally, the space dimension $N$ is assumed to be $N\geq 3$ in the zero mass case and $N\geq 1$ in the positive mass case.
+Finally, the space dimension NN is assumed to be $N\geq 3$ in the zero mass case and $N\geq 1$ in the positive mass case.
 We first prove that the functional $\Psi$ restricted on $D(\Psi)$ is continuous:
@@
 where $a\in(0,\infty)$ is any fixed number.
-Lemma 2.2.
+Lemma $2.$
 Suppose that $(u_{n})_{n\in\mathbb{N}}\subset D(\Psi)$ satisfies $u_{n}\rightharpoonup u$ weakly in $X$ and $\Psi_{0}(u_{n})\to\Psi_{0}(u)$. Then $\|u_{n}-u\|_{X}\to 0$ holds.
@@
 ∎
-Finally, we prove that every critical point of $I$ in $X$ becomes a classical solution of (1.1) and satisfies the Pohožaev identity.
+Finally, we prove that every critical point of II in $X$ becomes a classical solution of (1.1) and satisfies the Pohožaev identity.
 A similar assertion is obtained in [18, Propositions 2.6 and 2.10] for $N\geq 3$.
@@
 Lemma 2.3.
-Suppose either $N\geq 3$ and (F1)–(F2) or else $N\geq 1$ and (G1). Then each critical point $u$ of $I$ in $X$ is a classical solution of (1.1) and the Pohožaev identity $M(u)=0$ holds.
+Suppose either $N\geq 3$ and (F1)–(F2) or else $N\geq 1$ and (G1). Then each critical point $u$ of II in $X$ is a classical solution of (1.1) and the Pohožaev identity $M(u)=0$ holds.
 Proof.
-Let $u\in X$ be a critical point of $I$ and define
+Let $u\in X$ be a critical point of II and define
 $\rho(x):=\begin{dcases}f(u(x))&\text{(the zero mass case)},\\ -u(x)+g(u(x))&\text{(the positive mass case)}.\end{dcases}$
@@
 Hence, in this case, $u\in{\mathcal{C}}^{1,\alpha}(\mathbb{R}^{2})$ is a strong solution of (2.5) and the argument above leads to $u\in{\mathcal{C}}^{2}(\mathbb{R}^{2})$.
-Let $N=1$ and $u\in X$ be a critical point of $I$, and notice that $\rho\in L^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ in view of (G1) and (2.2).
+Let $N=1$ and $u\in X$ be a critical point of II, and notice that $\rho\in L^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ in view of (G1) and (2.2).
 We follow the argument in [8, Theorem 1.4, Proposition 2.7, Remark 2.9].
@@
 3. Concentration-compactness and a profile decomposition
-In this section, we establish a profile decomposition for bounded sequences in $X$, which is useful to prove the existence of minimizers of $I$ on ${\mathcal{M}}_{i}$.
+In this section, we establish a profile decomposition for bounded sequences in $X$, which is useful to prove the existence of minimizers of II on ${\mathcal{M}}_{i}$.
 Recall that in this section, the same notation to Section˜2 is used: $p_{0}$, $X$ and $\Psi_{0}$.
@@
 Finally we prove the profile decomposition.
-Proposition 3.3.
+Proposition $3.$
 Let $(u_{n})_{n\in\mathbb{N}}\subset X$ be bounded in $X$.
@@
 (4.3) $c_{0,MP}:=\inf_{\gamma\in\Gamma}\sup_{\tau\in[0,1]}I(\gamma(\tau)),\quad\Gamma:=\left\{\gamma\in{\mathcal{C}}([0,1];X_{0})\ :\ \gamma(0)=0,\ I(\gamma(1))<0\right\}.$
-Remark 4.4.
+Remark $4.$
 Fix any $\varphi\in{\mathcal{C}}^{\infty}_{0}(\mathbb{R}^{N})$ so that $\|\nabla\varphi\|_{\infty}<1$.
@@
 4.3. Minimizers on the Pohožaev set are critical points
-The aim of this subsection is to prove that every minimizer corresponding to $c_{0}$ is a critical point of $I$, and hence Section˜2 implies that it is a classical solution of (1.1).
+The aim of this subsection is to prove that every minimizer corresponding to $c_{0}$ is a critical point of II, and hence Section˜2 implies that it is a classical solution of (1.1).
 Lemma 4.6.
-Let $u_{0}\in{\mathcal{M}}_{0}$ satisfy $I(u_{0})=c_{0}$. Then $u_{0}$ is a critical point of $I$.
+Let $u_{0}\in{\mathcal{M}}_{0}$ satisfy $I(u_{0})=c_{0}$. Then $u_{0}$ is a critical point of II.
 Proof.
@@
 (4.4) $\frac{1}{2}\leq\theta(u)\leq 2,\quad\left|I(u_{\theta(u)})-I(u)\right|\leq\left|\int_{\theta(u)}^{1}\frac{d}{d\theta}I(u_{\theta})\,d\theta\right|\leq\frac{L_{1}}{2}|1-\theta(u)|^{2}.$
-To prove that $u_{0}$ is a critical point of $I$, let $v\in D(\Psi)$ be any point: otherwise, (1.5) holds trivially. For $t\in(0,1)$, we introduce
+To prove that $u_{0}$ is a critical point of II, let $v\in D(\Psi)$ be any point: otherwise, (1.5) holds trivially. For $t\in(0,1)$, we introduce
 $v_{t}:=(1-t)u_{0}+tv\in D(\Psi).$
@@
 $I(w_{n})-I(v_{t_{n}})\leq\frac{L_{1}}{2}|\theta_{n}-1|^{2}=O(t_{n}^{2})$
-and recalling that $u_{0}$ minimizes $I$ on ${\mathcal{M}}_{0}$, we see from (4.8) that
+and recalling that $u_{0}$ minimizes II on ${\mathcal{M}}_{0}$, we see from (4.8) that
 $0\leq I(w_{n})-I(u_{0})\leq I(v_{t_{n}})-I(u_{0})+O(t_{n}^{2})\leq t_{n}\left\{\Psi(v)-\Psi(u_{0})-\Phi^{\prime}(u_{0})(v-u_{0})\right\}+o(t_{n}).$
@@
 $0<\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx=\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx\leq C_{1}\limsup_{n\to\infty}\frac{t_{n}}{1-\theta_{n}^{N}}=0.$
-Thus (4.9) does not occur and $u_{0}$ is a critical point of $I$. ∎
-4.4. Existence of a minimizer and other properties of minimizers
+Thus (4.9) does not occur and $u_{0}$ is a critical point of II. ∎
+$4.$ Existence of a minimizer and other properties of minimizers
 To prove Theorem˜1.2, from Sections˜2, 4.2 and 4.3, it suffices to prove that there exists a minimizer, any minimizer has a constant sign and there exists a radial minimizer corresponding to $c_{0}$.
@@
 Furthermore, by $I(u_{n})\to c_{0}=I(\widetilde{u}_{j_{0}})$, we have $\Psi(u_{n})\to\Psi(\widetilde{u}_{j_{0}})$, and Section˜2 yields $\|u_{n}(\cdot+y_{n}^{j_{0}})-\widetilde{u}_{j_{0}}\|_{X_{0}}\to 0$.
-Thanks to Subsection˜4.3 we know that $\widetilde{u}_{j_{0}}$ is a critical point of $I$ and Section˜2 implies that $\widetilde{u}_{j_{0}}$ is a classical solution of (1.1).
+Thanks to Subsection˜4.3 we know that $\widetilde{u}_{j_{0}}$ is a critical point of II and Section˜2 implies that $\widetilde{u}_{j_{0}}$ is a classical solution of (1.1).
 We next prove that all minimizers have a constant sign.
@@
 (5.5) $\frac{d}{d\theta}\int_{\mathbb{R}^{N}}\frac{1}{2}g(\theta u)\theta u-G(\theta u)\,dx\geq 0.$
-To prove that minimizers on ${\mathcal{M}}_{1}$ are critical points of $I$, we need the following:
+To prove that minimizers on ${\mathcal{M}}_{1}$ are critical points of II, we need the following:
 Lemma 5.3.
@@
 Hence, $M_{1}(\gamma(1))<0$ and there exists $\tau\in(0,1)$ such that $M_{1}(\gamma(\tau_{1}))=0$. This completes the proof. ∎
-Lemma 5.5.
-Let $u_{0}\in{\mathcal{M}}_{1}$ satisfy $I(u_{0})=c_{1}$. Then $u_{0}$ is a critical point of $I$.
+Lemma $5.$
+Let $u_{0}\in{\mathcal{M}}_{1}$ satisfy $I(u_{0})=c_{1}$. Then $u_{0}$ is a critical point of II.
 Proof.
@@
 $=\frac{1}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx-\frac{(1-\theta_{n}^{N})}{N}\int_{\mathbb{R}^{N}}f(v_{t_{n}})v_{t_{n}}\,dx.$
-By $0<c_{1}=I(u_{0})=\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx/N$, we can repeat the same reasoning as for the zero mass case in Subsection˜4.3 to show that every minimizer of $I$ on ${\mathcal{M}}_{1}$ is a critical point.
+By $0<c_{1}=I(u_{0})=\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx/N$, we can repeat the same reasoning as for the zero mass case in Subsection˜4.3 to show that every minimizer of II on ${\mathcal{M}}_{1}$ is a critical point.
 ∎
@@
 $\int_{\mathbb{R}^{N}}g(u_{n})u_{n}-2G(u_{n})\,dx=\Phi^{\prime}(u_{n})u_{n}-2\Phi(u_{n})$
-is bounded and that each $u_{n}$ is $1$-Lipschitz in view of $u_{n}\in D(\Psi)$, we observe from (5.2) and (5.6) that $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{\infty}(\mathbb{R}^{N})$.
+is bounded and that each $u_{n}$ is 11-Lipschitz in view of $u_{n}\in D(\Psi)$, we observe from (5.2) and (5.6) that $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{\infty}(\mathbb{R}^{N})$.
 Furthermore, by (5.2) and (5.6), for each $t>0$ there exists $\delta_{t}>0$ such that $g(s)s-2G(s)\geq\delta_{t}$ if $|s|\geq t$.
-Therefore, $(|\{|u_{n}|\geq t\}|)_{n\in\mathbb{N}}$ is also bounded where $|A|$ denotes the $N$-dimensional Lebesgue measure of $A\subset\mathbb{R}^{N}$.
+Therefore, $(|\{|u_{n}|\geq t\}|)_{n\in\mathbb{N}}$ is also bounded where $|A|$ denotes the NN-dimensional Lebesgue measure of $A\subset\mathbb{R}^{N}$.
 By (G1), choose $t_{0}>0$ so that $|G(t)|\leq t^{2}/4$ for each $|t|\leq t_{0}$.
@@
 $=\liminf_{n\to\infty}I(u_{n})=c_{1}.$
-Therefore, $\theta_{j_{0}}=1$ and $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{1}$ is a minimizer. Furthermore, by Subsection˜5.2, $\widetilde{u}_{j_{0}}$ is a critical point of $I$.
+Therefore, $\theta_{j_{0}}=1$ and $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{1}$ is a minimizer. Furthermore, by Subsection˜5.2, $\widetilde{u}_{j_{0}}$ is a critical point of II.
 We next prove $\|u_{n}(\cdot+y^{j_{0}}_{n})-\widetilde{u}_{j_{0}}\|_{X_{1}}\to 0$.
@@
 $g$) is odd, $\Psi(-u)=\Psi(u)$ and $\Phi(-u)=\Phi(u)$.
-According to the principle of symmetric criticality (see [27, 18]), any critical point of $I$ in $X_{\tau}$ is also a critical point in $X_{\mathcal{O}}$, and by applying the principle again, it is a critical point in $X$.
+According to the principle of symmetric criticality (see [27, 18]), any critical point of II in $X_{\tau}$ is also a critical point in $X_{\mathcal{O}}$, and by applying the principle again, it is a critical point in $X$.
 Thus, it suffices to find a critical point of $I|_{X_{\tau}}$.
@@
 Let $u\in{\mathcal{M}}_{\tau}$ satisfy $I(u)=\inf_{{\mathcal{M}}_{\tau}}I$.
-Then we may use the argument in the proof of Subsections˜4.3 and 5.2 to verify that $u$ is a critical point of $I$ on $X_{\tau}$.
-As pointed in the beginning of this section, $u$ is a critical point of $I$ in $X$, and hence Section˜2 implies that $u$ is a classical solution of (1.1).
+Then we may use the argument in the proof of Subsections˜4.3 and 5.2 to verify that $u$ is a critical point of II on $X_{\tau}$.
+As pointed in the beginning of this section, $u$ is a critical point of II in $X$, and hence Section˜2 implies that $u$ is a classical solution of (1.1).
 When $k_{1}=k_{2}$, we next prove $I(u)\geq 2\inf_{{\mathcal{M}}}I$ where $u\in{\mathcal{M}}_{\tau}$ is a minimizer corresponding to $\inf_{{\mathcal{M}}_{\tau}}I$. Set